### v2

# Using Dynamic Safety Margins as Control Barrier Functions

Abstract
This paper provides a systematic approach to design control barrier functions (CBFs) for arbitrary state and input constraints using tools from the reference governor literature. In particular, it is shown that dynamic safety margins (DSMs) are CBFs for an augmented system obtained by concatenating the state and a virtual reference. The proposed approach is agnostic to the relative degree and can handle multiple state and input constraints using the control-sharing property of CBFs. The construction of CBFs using Lyapunov-based DSMs is then investigated in further detail. Numerical simulations show that the method outperforms existing DSM-based approaches, while also guaranteeing safety and persistent feasibility of the associated optimization program.
{IEEEkeywords}
Constrained control, Control barrier functions, Predictive control for nonlinear systems.
1 Introduction
[…]
Control barrier functions (CBFs) have become a popular tool for deriving optimization-based constrained control laws that are safe, easy to implement, and achieve good performance [1, 2, 3]. Despite their success, the widespread use of CBFs is limited by the absence of a systematic method to synthesize them for general classes of systems with arbitrary state and input constraints.
Given a smooth state constraint set, it is usually straightforward to design a differentiable function of the state vector that is positive for constraint-admissible states, and negative otherwise. We call such functions candidate CBFs. They are commonly used in practice to design CBF-based controllers and, with tuning, can achieve great performance (e.g., [2]). However, candidate CBFs lack many safety guarantees that valid CBFs provide. In particular, the CBF-based optimization program governing the control law may become infeasible when using candidate CBFs, which is unacceptable for safety-critical systems. Ultimately, CBFs are certificates of control invariance. Therefore, if the original state constraint set is not control invariant, it is impossible to find a valid CBF associated to it. In these cases, we must resort to finding CBFs for a, hopefully large, control invariant subset of the state constraint set. In this work, we use tools from the reference governor literature [4] to systematically design valid CBFs given arbitrary state and input constraints.
Figure 1: Comparison of CBF-based control (top), which filters a nominal input signal $\boldsymbol{\mathbf{u}}_{n}$ to obtain a safe input $\boldsymbol{\mathbf{u}}^{*}$, with a specific reference governor approach, the command governor (CG). The CG filters the desired reference signal $\boldsymbol{\mathbf{r}}$ to obtain a safe virtual reference $\boldsymbol{\mathbf{v}}^{*}$ for the prestabilizing controller.
1.1 Related Works
While the term control barrier function can be traced back to the early 2000’s (see [5] for a historical account), the modern definition that we consider in this work was first presented in [6] under the name of zeroing CBFs. Since then, many works have studied different properties of CBFs and the systems they are used in: [7, 8] address the relative degree of CBFs; [9] defines control-sharing CBFs to enforce multiple constraints at once; [10, 11] consider discrete-time and sampled-data systems, respectively; [12, 13] study robust CBFs.
Most relevant to the present manuscript, we review some works that study the design of CBFs and feasibility of the CBF-based optimization program. In [14], sum of squares (SOS) techniques are used to synthesize CBFs for semi-algebraic sets and polynomial systems using an alternating descent algorithm. In [15], the authors use SOS techniques to synthesize polynomial control Lyapunov functions, and then present an extension to construct CBFs. In [16], the authors introduce a variable modifying the decay rate of the CBF to guarantee feasibility of the CBF-based program when the state is in the interior of the safe set even under input constraints. However, their approach provides no feasibility guarantee when the state is on the boundary of the safe set. In [17], the authors design CBFs for Euler–Lagrange systems. While the results are promising and the class of systems is relevant, the approach is limited to box constraints. In [18], the authors learn CBFs from human demonstrations but they provide no guarantees that the learned CBFs are valid. In [19], given a valid CBF, the authors use learning to construct a CBF for a larger control invariant set using a combination of safe and unsafe trajectories of the system. Nonetheless, finding the initial, valid CBF remains an open question. In [20], the authors use a backup control policy to enlarge a small (but easy to find) control invariant set. The enlarged control invariant set is then used to derive a backup CBF. These backup CBFs rarely have a closed form, which makes them difficult to implement. In [21], the authors use maximal output admissible sets (MOASs) to design discrete-time CBFs for arbitrary state and input constraints. While the approach works well for linear systems, finding the MOAS for nonlinear systems remains an open question.
This work expands the underlying theme of [20, 21], where CBFs are obtained starting from a prestabilizing (or backup) controller. To do this, we adopt the notion of dynamic safety margin (DSM) from the explicit reference governor (ERG) framework [22]. Given a fixed reference, a DSM measures the distance to constraint violation of a prestabilized system trajectory as it converges to the associated equilibrium point. In [23], the authors define governor-parameterized barrier functions (PBFs), which are based on CBFs, and show that they are DSMs. While the results are limited to linear systems, their work is one of the few connections between the reference governor and CBF literatures. In contrast to their result, we show that DSMs are CBFs for an augmented system consisting of the concatenated state and prestabilizing controller reference vectors. These works serve as motivation for exploring the connections between reference governor approaches like the command governor (CG) [24], which filters a nominal reference, and CBF-based control, which filters a nominal input. Both approaches ultimately modify a nominal, but potentially unsafe signal to guarantee closed-loop safety (see Fig. 1).
1.2 Contributions
The contributions of this manuscript are:
  • We show that dynamic safety margins (DSMs) are control barrier functions (CBFs) for an augmented system consisting of the concatenation of the state vector and the reference of the prestabilizing controller used to construct the DSM.
  • We show that the DSM-based CBFs support constraints with arbitrary relative degree.
  • We show that DSMs obtained for the same prestabilizing controller are control-sharing CBFs. That is, they can be simultaneously imposed in the CBF-based program and feasibility is retained.
  • We extend the Lyapunov-based DSMs presented in [22] to inadmissible references. This extension is needed for DSM-based CBFs.
  • We provide conditions on the DSM and the constraints that guarantee the CBF-based controller is locally Lipschitz contionuous. This guarantees uniqueness of the closed-loop trajectory.
  • We present a comparison of our proposed approach with the explicit reference governor (ERG) [22] and candidate CBFs in two nonlinear systems. The numerical simulations show that the DSM-based CBFs we propose outperform the ERG, while boasting strict safety guarantees that are otherwise lost when candidate CBF-based control becomes infeasible.
The rest of the paper is organized as follows: Section 2 provides formal definitions for control invariance, CBFs, and DSMs; Section 3 proves that DSMs are CBFs, shows the relative degree of the constructed CBF is irrelevant for safety, and formulates the CBF-based optimization problem, which is a quadratic program (QP) for polyhedral input constraints; Section 4 presents the control-sharing property of CBFs and shows that DSMs have it; Section 5 provides a systematic approach based on Lyapunov functions to construct DSMs and highlights some desirable properties that such DSMs have; Section 6 presents the numerical simulations and comparisons; Section 7 concludes the paper.
1.3 Notation
Throughout the paper, only vector quantities are presented in bold font: $\boldsymbol{\mathbf{x}}\in\mathbb{R}^{n}$. The notation $[\boldsymbol{\mathbf{x}};~{}\boldsymbol{\mathbf{y}}]$ denotes the vertical stacking of two vectors $\boldsymbol{\mathbf{x}}$ and $\boldsymbol{\mathbf{y}}$ into a column vector. A continuous function $\alpha:[0,\infty)\to[0,\infty)$ is said to belong to class $\mathscr{K}_{\infty}$ if it is strictly increasing, $\alpha(0)=0$, and $\lim_{r\to\infty}\alpha(r)=\infty$. The symbol $\subset$ denotes set inclusion, but not necessarily strict inclusion. Given a set $\mathcal{S}$, we denote by $\operatorname{Int}\left(\mathcal{S}\right)$ the interior of $\mathcal{S}$, by ${\overline{\mathcal{S}}}$ the closure of $\mathcal{S}$ and by $\partial\mathcal{S}$ the boundary of $\mathcal{S}$ (i.e. $\partial\mathcal{S}={\overline{\mathcal{S}}}\setminus\operatorname{Int}\left(\mathcal{S}\right)$). Let $h:\mathcal{D}\to\mathbb{R}$ be continuously differentiable. We say $\boldsymbol{\mathbf{x}}\in\mathcal{D}$ is a regular point of $h$ if $\partial h/\partial\boldsymbol{\mathbf{x}}\neq 0$. We say $c\in\mathbb{R}$ is a regular value of $h$ if every $\boldsymbol{\mathbf{x}}\in\mathcal{D}$ such that $h(\boldsymbol{\mathbf{x}})=c$ is a regular point.
2 Preliminaries
This section recalls the notion of control invariance, control barrier functions, and dynamic safety margins. Let $\mathbb{X}\subset\mathbb{R}^{n}$ and $\mathbb{U}\subset\mathbb{R}^{m}$ be the state and input spaces, respectively, and consider a control-affine system
$\dot{\boldsymbol{\mathbf{x}}}=f(\boldsymbol{\mathbf{x}})+g(\boldsymbol{\mathbf{x}})\boldsymbol{\mathbf{u}},$ (1)
where $f:\mathbb{X}\to\mathbb{R}^{n}$ and $g:\mathbb{X}\to\mathbb{R}^{n\times m}$ are locally Lipschitz continuous functions. The system is subject to state $\boldsymbol{\mathbf{x}}\in\mathcal{X}$ and input $\boldsymbol{\mathbf{u}}\in\mathcal{U}$ constraints, where $\mathcal{X}\subset\mathbb{X}$ and $\mathcal{U}\subset\mathbb{U}$ are closed sets.
2.1 Control Invariance
Control invariance is a desirable property of state-space subsets that is closely related to safety. Roughly, a set is control invariant if there exists a suitable choice of inputs such that the trajectory of the closed-loop system remains in the set.
Definition 1
[25] A compact set $\mathcal{C}\subset\mathbb{X}$ is control invariant if, $\forall\boldsymbol{\mathbf{x}}_{0}\in\mathcal{C}$, there exists an admissible input signal $\boldsymbol{\mathbf{u}}:[0,\infty)\to\mathcal{U}$ such that, if they exist, at least one of the trajectories of the closed-loop system (1), with $\boldsymbol{\mathbf{x}}(0)=\boldsymbol{\mathbf{x}}_{0}$, satisfies $\boldsymbol{\mathbf{x}}(t)\in\mathcal{C}$ for all $t$ in its interval of existence $\mathcal{J}(\boldsymbol{\mathbf{x}}_{0})\subset[0,\infty)$.
Note that this concept is sometimes called weak control invariance [25] because we don’t guarantee the existence and uniqueness of the trajectory. Nonetheless, it will simplify the presentation of the upcoming results. Assumptions for the existence and uniqueness of solutions will be presented separately. Central to the upcoming discussion is Nagumo’s Theorem characterizing control invariance in terms of Bouligand’s tangent cone.
Theorem 1 (Nagumo)
[26] A compact set $\mathcal{C}\subset\mathbb{X}$ is control invariant if and only if
$\forall\boldsymbol{\mathbf{x}}\in\mathcal{C},\quad\exists\boldsymbol{\mathbf{u}}\in\mathcal{U},\quad f(\boldsymbol{\mathbf{x}})+g(\boldsymbol{\mathbf{x}})\boldsymbol{\mathbf{u}}\in\mathcal{T}_{\mathcal{C}}(\boldsymbol{\mathbf{x}}),$ (2)
where $\mathcal{T}_{\mathcal{C}}(\boldsymbol{\mathbf{x}})$ is the tangent cone to $\mathcal{C}$ in $\boldsymbol{\mathbf{x}}$ [25, Definition 3.1].
While this characterization was groundbreaking, working with the tangent cone $\mathcal{T}_{\mathcal{C}}(\boldsymbol{\mathbf{x}})$ is cumbersome. Fortunately, when the compact set $\mathcal{C}$ is smooth, the tangent cone in the boundary is reduced to the tangent half-space shifted to the origin and we have the following characterization.
Theorem 2
[27, Proposition 4.3.7] Let $h:\mathcal{D}\to\mathbb{R}$ be a continuously differentiable function with regular value 00 and suppose $\mathcal{C}=\{\boldsymbol{\mathbf{x}}\in\mathcal{D}:h(\boldsymbol{\mathbf{x}})\geq 0\}\subset\mathbb{X}$ is compact. The set $\mathcal{C}$ is control invariant if and only if
$\forall\boldsymbol{\mathbf{x}}\in\partial\mathcal{C},\quad\exists\boldsymbol{\mathbf{u}}\in\mathcal{U},\quad L_{f}h(\boldsymbol{\mathbf{x}})+L_{g}h(\boldsymbol{\mathbf{x}})\boldsymbol{\mathbf{u}}\geq 0,$ (3)
where $L_{f}h$ and $L_{g}h$ are the Lie derivatives of $h$ along $f$ and $g$, respectively.
Although much simpler than evaluating the tangent cone, the conditions of Theorem 2 apply only on the boundary of the set $\partial\mathcal{C}$, which makes implementing them difficult in practice. This motivated the development of control barrier functions, which allow us to extend these conditions to the entirety of the set $\mathcal{C}$.
2.2 Control Barrier Functions
Control barrier functions are real-valued functions that certify the control invariance of their superlevel set.
Definition 2
[6] A continuously differentiable function $h:\mathcal{D}\to\mathbb{R}$ is a control barrier function (CBF) if there exists a class $\mathscr{K}_{\infty}$ function $\alpha:[0,\infty)\to[0,\infty)$ such that
$\forall\boldsymbol{\mathbf{x}}\in\mathcal{C},\quad\sup_{\boldsymbol{\mathbf{u}}\in\mathcal{U}}\left[L_{f}h(\boldsymbol{\mathbf{x}})+L_{g}h(\boldsymbol{\mathbf{x}})\boldsymbol{\mathbf{u}}\right]\geq-\alpha\big{(}h(\boldsymbol{\mathbf{x}})\big{)},$ (4)
where
$\mathcal{C}=\{\boldsymbol{\mathbf{x}}\in\mathcal{D}\mid h(\boldsymbol{\mathbf{x}})\geq 0\}\subset\mathbb{X}.$ (5)
With this definition, the conditions of Theorem 2 were extended to the entire set $\mathcal{C}$, as shown in the next result.
Lemma 1
Assume $\mathcal{U}\subset\mathbb{U}$ is compact. Let $\mathcal{C}\subset\mathbb{X}$, defined as in (5), be compact. Moreover, let 00 be a regular value of $h$. The set $\mathcal{C}$ is control invariant if and only if $h$ is a CBF.
Proof 2.3.
Although the proof is based on [6], they showed this result for barrier functions, not control barrier functions. Let $\mathcal{C}\subset\mathbb{X}$ be control invariant and define the set $\mathcal{C}_{c}=\{\boldsymbol{\mathbf{x}}\in\mathcal{D}:0\leq h(\boldsymbol{\mathbf{x}})\leq c\}$. Note that for any $c\geq 0$, $\mathcal{C}_{c}\subset\mathcal{C}$ is compact. Next, define the function $\hat{\alpha}:[0,\infty)\to\mathbb{R}$
$\hat{\alpha}(c)=-\inf_{\boldsymbol{\mathbf{x}}\in\mathcal{C}_{c}}\sup_{\boldsymbol{\mathbf{u}}\in\mathcal{U}}\left[L_{f}h(\boldsymbol{\mathbf{x}})+L_{g}h(\boldsymbol{\mathbf{x}})\boldsymbol{\mathbf{u}}\right].$ (6)
By compactness of $\mathcal{U}$, the function $\xi:\mathcal{D}\to\mathbb{R}$ defined as
$\xi(\boldsymbol{\mathbf{x}})=\sup_{\boldsymbol{\mathbf{u}}\in\mathcal{U}}\left[L_{f}h(\boldsymbol{\mathbf{x}})+L_{g}h(\boldsymbol{\mathbf{x}})\boldsymbol{\mathbf{u}}\right],$ (7)
is continuous on $\mathcal{D}$, making $\hat{\alpha}$ also continuous and non-decreasing. Furthermore, $\forall\boldsymbol{\mathbf{x}}\in\mathcal{C}$, $\exists\boldsymbol{\mathbf{u}}\in\mathcal{U}$, such that $L_{f}h(\boldsymbol{\mathbf{x}})+L_{g}h(\boldsymbol{\mathbf{x}})\boldsymbol{\mathbf{u}}\geq-\hat{\alpha}\big{(}h(\boldsymbol{\mathbf{x}})\big{)}$. To see why, let $\boldsymbol{\mathbf{x}}_{0}\in\mathcal{C}$ be given and define $c_{0}=h(\boldsymbol{\mathbf{x}}_{0})$. There exists $\boldsymbol{\mathbf{u}}_{0}\in\mathcal{U}$ such that
$L_{f}h(\boldsymbol{\mathbf{x}}_{0})+L_{g}h(\boldsymbol{\mathbf{x}}_{0})\boldsymbol{\mathbf{u}}_{0}=\xi(\boldsymbol{\mathbf{x}}_{0})\geq\inf_{\boldsymbol{\mathbf{x}}\in\mathcal{C}_{c_{0}}}\xi(\boldsymbol{\mathbf{x}})=-\hat{\alpha}\big{(}h(\boldsymbol{\mathbf{x}}_{0})\big{)}$
By Theorem 2, for any $\boldsymbol{\mathbf{x}}\in\partial\mathcal{C}$, $\exists\boldsymbol{\mathbf{u}}\in\mathcal{U}$ such that $L_{f}h(\boldsymbol{\mathbf{x}})+L_{g}h(\boldsymbol{\mathbf{x}})\boldsymbol{\mathbf{u}}\geq 0$. Thus, $\forall\boldsymbol{\mathbf{x}}\in\partial\mathcal{C}=\mathcal{C}_{0}$, $\xi(\boldsymbol{\mathbf{x}})\geq 0$. As a result,
$\hat{\alpha}(0)=-\inf_{\boldsymbol{\mathbf{x}}\in\partial\mathcal{C}}\xi(\boldsymbol{\mathbf{x}})\leq 0$ (8)
There always exists a class $\mathscr{K}_{\infty}$ function $\alpha:[0,\infty)\to[0,\infty)$ that upper-bounds $\hat{\alpha}$, yielding that $\forall\boldsymbol{\mathbf{x}}\in\mathcal{C}$,
$\sup_{\boldsymbol{\mathbf{u}}\in\mathcal{U}}\left[L_{f}h(\boldsymbol{\mathbf{x}})+L_{g}h(\boldsymbol{\mathbf{x}})\boldsymbol{\mathbf{u}}\right]\geq-\hat{\alpha}\big{(}h(\boldsymbol{\mathbf{x}})\big{)}\geq-\alpha\big{(}h(\boldsymbol{\mathbf{x}})\big{)}.$
On the other hand, assume $h$ is a CBF, let $\alpha\in\mathscr{K}_{\infty}$ be such that (4) holds, and let $\boldsymbol{\mathbf{x}}\in\partial\mathcal{C}$ be given. By compactness of $\mathcal{U}$, there exists $\boldsymbol{\mathbf{u}}\in\mathcal{U}$ such that $L_{f}h(\boldsymbol{\mathbf{x}})+L_{g}h(\boldsymbol{\mathbf{x}})\boldsymbol{\mathbf{u}}\geq-\alpha\big{(}h(\boldsymbol{\mathbf{x}})\big{)}=0$, where the last equality holds because $h(\boldsymbol{\mathbf{x}})=0$. Since $\mathcal{C}$ is compact and 00 is a regular value of $h$, we can apply Theorem 2 to conclude $\mathcal{C}$ is control invariant.
Notice compactness of $\mathcal{U}$ is necessary for continuity of the function $\xi:\mathcal{D}\to\mathbb{R}$, and therefore of $\hat{\alpha}:[0,\infty)\to\mathbb{R}$. Without it, we lose necessity in Lemma 1. However, the sufficiency in Lemma 1 (i.e., the CBF certifies control invariance of its compact superlevel set) is maintained even when the compactness of $\mathcal{U}$ is exchanged for the less restrictive condition
$\forall\boldsymbol{\mathbf{x}}\in\mathcal{C},\quad\mathcal{K}(\boldsymbol{\mathbf{x}})\neq\emptyset,$ (9)
where
$\mathcal{K}(\boldsymbol{\mathbf{x}})=\{\boldsymbol{\mathbf{u}}\in\mathcal{U}\mid L_{f}h(\boldsymbol{\mathbf{x}})+L_{g}h(\boldsymbol{\mathbf{x}})\boldsymbol{\mathbf{u}}+\alpha\big{(}h(\boldsymbol{\mathbf{x}})\big{)}\geq 0\}.$ (10)
In this work we drop the compactness requirement on $\mathcal{U}$ because only the sufficiency statement in Lemma 1 is relevant and we will design CBFs for which (9) holds.
CBFs can be used to design add-on modules that enforce constraint satisfaction by filtering the control input. Specifically, let $h$ be a CBF such that $\mathcal{C}\subset\mathcal{X}$ and consider the closed-loop system
$\dot{\boldsymbol{\mathbf{x}}}=f(\boldsymbol{\mathbf{x}})+g(\boldsymbol{\mathbf{x}})\beta(\boldsymbol{\mathbf{x}}),$ (11)
where $\beta:\mathcal{C}\to\mathcal{U}$ is the optimization-based control policy
$\beta(\boldsymbol{\mathbf{x}})=\underset{\boldsymbol{\mathbf{u}}\in\mathcal{K}(\boldsymbol{\mathbf{x}})}{\operatorname{argmin}}\>\|\boldsymbol{\mathbf{u}}-\kappa(\boldsymbol{\mathbf{x}})\|^{2},$ (12)
and $\kappa:\mathbb{X}\to\mathbb{U}$ is a nominal control policy with desirable closed-loop performance. When (9) holds, the optimization problem (12) is always feasible, making $\beta$ well-defined. If, in addition, the policy $\beta$ is locally Lipschitz, we can guarantee the existence and uniqueness of a solution $\boldsymbol{\mathbf{x}}(t)$ to the closed-loop system (11) and that it will satisfy $\boldsymbol{\mathbf{x}}(t)\in\mathcal{X}$. In upcoming sections, we will specify conditions under which $\beta$ is locally Lipschitz.
As noted in the introduction, finding a CBF for arbitrary constraint sets $\mathcal{X}\subset\mathbb{X}$ is a challenging task. In fact, Lemma 1 states that, when $\mathcal{U}$ is compact, a CBF $h$ with a well-defined policy $\beta$ exists only if $\mathcal{X}$ is control invariant, which is generally not the case. In spite of this, many practitioners insist on imposing $\mathcal{C}=\mathcal{X}$ to construct a “candidate” CBF (i.e. a continuously differentiable function $h:\mathcal{D}\to\mathbb{R}$ that satisfies $\mathcal{X}=\{\boldsymbol{\mathbf{x}}\in\mathcal{D}:h(\boldsymbol{\mathbf{x}})\geq 0\}$) and then tuning a class $\mathscr{K}_{\infty}$ function $\alpha$ on a case-by-case basis. Although this approach can yield good performance, it suffers from the possibility that $\mathcal{K}(\boldsymbol{\mathbf{x}})=\emptyset$ for some $x\in\mathcal{X}$, which makes the controller $\beta$ ill-defined.
In this paper, we leverage results from the ERG framework to design a valid CBF that satisfies Definition 2 and, additionally, is guaranteed to satisfy (9).
2.3 Dynamic Safety Margins
Dynamic safety margins are real-valued functions that quantify the distance to constraint violation for a prestabilized system subject to a constant reference. In order to use them, we must make the following assumptions about system (1).
Assumption 1
System (1) admits a simply connected equilibrium manifold parameterized by continuous functions $\bar{x}:\mathbb{V}\to\mathbb{X}$ and $\bar{u}:\mathbb{V}\to\mathbb{U}$ such that
$\forall\boldsymbol{\mathbf{v}}\in\mathbb{V},\quad f\big{(}\bar{x}(\boldsymbol{\mathbf{v}})\big{)}+g\big{(}\bar{x}(\boldsymbol{\mathbf{v}})\big{)}\bar{u}(\boldsymbol{\mathbf{v}})=0,$ (13)
where $\mathbb{V}\subset\mathbb{R}^{l}$ is the reference space. Moreover, every equilibrium point $\bar{x}(\boldsymbol{\mathbf{v}})$ is stabilizable.
The vector $\boldsymbol{\mathbf{v}}\in\mathbb{V}$ is hereafter referred to as a “reference” for the system because it identifies a specific equilibrium point. The set of steady-state admissible references is
$\mathcal{V}=\{\boldsymbol{\mathbf{v}}\in\mathbb{V}\mid\bar{x}(\boldsymbol{\mathbf{v}})\in\mathcal{X},~{}\bar{u}(\boldsymbol{\mathbf{v}})\in\mathcal{U}\}.$ (14)
Note that, since $\bar{x}$, $\bar{u}$ are continuous and $\mathcal{X}$, $\mathcal{U}$ are closed sets, $\mathcal{V}$ is closed in $\mathbb{V}$.
Assumption 2
The set $\mathcal{V}\subset\mathbb{V}$ is not empty, i.e., system (1) features at least one steady-state admissible equilibrium point.
In essence, Assumption 1 states that the system has an open set of stabilizable equilibria, and Assumption 2 states that at least one of the equilibria is safe. By Assumption 1, let $\pi:\mathbb{X}\times\mathbb{V}\to\mathbb{U}$, locally Lipschitz continuous on $\mathbb{X}$ and continuous on $\mathbb{V}$, be a prestabilizing control law such that $\pi\big{(}\bar{x}(\boldsymbol{\mathbf{v}}),\boldsymbol{\mathbf{v}}\big{)}=\bar{u}(\boldsymbol{\mathbf{v}})$, and $\bar{x}(\boldsymbol{\mathbf{v}})$ is an asymptotically stable equilibrium point of the prestabilized system
$\dot{\boldsymbol{\mathbf{x}}}=f(\boldsymbol{\mathbf{x}})+g(\boldsymbol{\mathbf{x}})\pi(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\triangleq f_{\pi}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}}).$ (15)
Given a reference $\boldsymbol{\mathbf{v}}\in\mathbb{V}$, the prestabilized dynamics $f_{\pi}$ ensure the existence of an open set $\mathcal{D}_{\boldsymbol{\mathbf{v}}}\subset\mathbb{X}$ that satisfies $\bar{x}(\boldsymbol{\mathbf{v}})\in\mathcal{D}_{\boldsymbol{\mathbf{v}}}$ and is such that
$\boldsymbol{\mathbf{x}}(0)\in\mathcal{D}_{\boldsymbol{\mathbf{v}}}\quad\implies\quad\lim_{t\to\infty}\boldsymbol{\mathbf{x}}(t)=\bar{x}(\boldsymbol{\mathbf{v}}).$ (16)
The set $\mathcal{D}_{\boldsymbol{\mathbf{v}}}$ is a region of attraction for $\bar{x}(\boldsymbol{\mathbf{v}})$. We assume small changes in $\boldsymbol{\mathbf{v}}$ yield small changes in $\mathcal{D}_{\boldsymbol{\mathbf{v}}}\subset\mathbb{X}$. Formally,
Assumption 3
The set
$\mathcal{D}_{\boldsymbol{\mathbf{v}}}\times\mathbb{V}\triangleq\{(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\mathbb{X}\times\mathbb{V}\mid\boldsymbol{\mathbf{x}}\in\mathcal{D}_{\boldsymbol{\mathbf{v}}}\},$ (17)
is open in $\mathbb{X}\times\mathbb{V}$.
Note that, for each reference $\boldsymbol{\mathbf{v}}\in\mathbb{V}$, the input constraint set $\mathcal{U}$ induces state constraints on the prestabilized system. This allows us to define the reference-dependent state constraint set
$\mathcal{X}_{\boldsymbol{\mathbf{v}}}=\{\boldsymbol{\mathbf{x}}\in\mathcal{X}\mid\pi(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\mathcal{U}\}.$ (18)
By continuity of $\pi$ on $\mathbb{V}$, it follows that $\mathcal{X}_{\boldsymbol{\mathbf{v}}}$ is closed in $\mathbb{X}$. The following DSM definition satisfies the one proposed in [22, Definition 1], although it makes stronger assumptions by requiring a) differentiability rather than continuity and b) invariance rather than strong returnability.
Definition 2.4.
A continuously differentiable function $\Delta:\!\mathcal{D}_{\boldsymbol{\mathbf{v}}}\!\times\mathbb{V}\!\!\to\!\!\mathbb{R}$ is a dynamic safety margin (DSM) for $\pi$ if, $\forall\boldsymbol{\mathbf{v}}\!\in\!\mathbb{V}$,
$\boldsymbol{\mathbf{v}}\in\mathcal{V}~{}$ $\implies~{}\Delta\big{(}\bar{x}(\boldsymbol{\mathbf{v}}),\boldsymbol{\mathbf{v}}\big{)}\geq 0,$ (19a)
$\Delta(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\geq 0~{}$ $\implies~{}\boldsymbol{\mathbf{x}}\in\mathcal{X}_{\boldsymbol{\mathbf{v}}},$ (19b)
$\Delta(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\geq 0~{}$ $\implies~{}\frac{\partial\Delta}{\partial\boldsymbol{\mathbf{x}}}f_{\pi}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\geq 0,$ (19c)
$\epsilon>0~{}$ $\implies~{}\mathcal{C}_{\boldsymbol{\mathbf{v}}}^{\epsilon}~{}\text{is closed in $\mathbb{X}$,}$ (19d)
where $\mathcal{C}_{\boldsymbol{\mathbf{v}}}^{\epsilon}=\{\boldsymbol{\mathbf{x}}\in\mathcal{D}_{\boldsymbol{\mathbf{v}}}\mid\Delta(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\geq\epsilon\}$.
Lemma 2.5.
If $\Delta(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\geq 0$, then $\boldsymbol{\mathbf{v}}\in\mathcal{V}$.
Proof 2.6.
Let $(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\mathcal{D}_{\boldsymbol{\mathbf{v}}}\times\mathbb{V}$ be such that $\Delta(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{r}})\geq 0$. Given the initial condition $\boldsymbol{\mathbf{x}}(0)=\boldsymbol{\mathbf{x}}$, let $\dot{\boldsymbol{\mathbf{x}}}(t)=f_{\pi}\big{(}\boldsymbol{\mathbf{x}}(t),\boldsymbol{\mathbf{v}}\big{)}$. Then, it follows from the continuity of $\Delta$ and property (19c) that $\Delta\big{(}\boldsymbol{\mathbf{x}}(t),\boldsymbol{\mathbf{v}}\big{)}\geq 0$. Therefore, by property (19b), $\boldsymbol{\mathbf{x}}(t)\in\mathcal{X}_{\boldsymbol{\mathbf{v}}}$ for all $t\geq 0$. Taking the limit as $t\to\infty$, we obtain
$\lim_{t\to\infty}\boldsymbol{\mathbf{x}}(t)=\bar{x}(\boldsymbol{\mathbf{v}})\in\mathcal{X}_{\boldsymbol{\mathbf{v}}}\subset\mathcal{X}.$ (20)
Finally, $\pi\big{(}\bar{x}(\boldsymbol{\mathbf{v}}),\boldsymbol{\mathbf{v}}\big{)}=\bar{u}(\boldsymbol{\mathbf{v}})\in\mathcal{U}$, and $\boldsymbol{\mathbf{v}}\in\mathcal{V}$.
DSMs are used to design add-on modules (ERGs) that enforce constraint satisfaction by filtering the reference of the prestabilized system. Specifically, let us augment the prestabilized system with reference dynamics to obtain
$\begin{bmatrix}\dot{\boldsymbol{\mathbf{x}}}\\ \dot{\boldsymbol{\mathbf{v}}}\end{bmatrix}=\begin{bmatrix}f(\boldsymbol{\mathbf{x}})+g(\boldsymbol{\mathbf{x}})\pi(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\\ \Delta(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\rho(\boldsymbol{\mathbf{v}},\boldsymbol{\mathbf{r}})\end{bmatrix},$ (21)
where $\rho:\mathbb{V}\times\mathbb{V}\to\mathbb{R}^{l}$ is any locally Lipschitz continuous function such that $\boldsymbol{\mathbf{r}}\in\mathbb{V}$ is an asymptotically stable equilibrium point of the system $\dot{\boldsymbol{\mathbf{v}}}(t)=\rho\big{(}\boldsymbol{\mathbf{v}}(t),\boldsymbol{\mathbf{r}}\big{)}$. In the ERG literature, $\rho$ is called a navigation field [22, Definition 2]. System (21) ensures that, if $\big{(}\boldsymbol{\mathbf{x}}(0),\boldsymbol{\mathbf{v}}(0)\big{)}\in\mathcal{X}_{\boldsymbol{\mathbf{v}}(0)}\times\mathcal{V}$, then $\big{(}\boldsymbol{\mathbf{x}}(t),\boldsymbol{\mathbf{v}}(t)\big{)}\in\mathcal{X}_{\boldsymbol{\mathbf{v}}(t)}\times\mathcal{V}$ at all times [22]. The ERG literature provides systematic tools for computing DSMs for various classes of systems [28, 22, 29, 30, 31]. The following section shows how these tools from the ERG literature can be used to systematically construct valid CBFs.
3 From Dynamic Safety Margins to Control Barrier Functions
Based on the intuition that both CBFs and DSMs can be used to design add-on modules for constraint handling, the following theorem states that a dynamic safety margin is, in fact, a CBF for an augmented system consisting of the concatenation of $\boldsymbol{\mathbf{x}}\in\mathbb{X}\subset\mathbb{R}^{n}$ and $\boldsymbol{\mathbf{v}}\in\mathbb{V}\subset\mathbb{R}^{l}$
$\begin{bmatrix}\dot{\boldsymbol{\mathbf{x}}}\\ \dot{\boldsymbol{\mathbf{v}}}\end{bmatrix}=\begin{bmatrix}f(\boldsymbol{\mathbf{x}})+g(\boldsymbol{\mathbf{x}})\boldsymbol{\mathbf{u}}\\ \boldsymbol{\mathbf{w}}\end{bmatrix},$ (22)
where $(\boldsymbol{\mathbf{u}},\boldsymbol{\mathbf{w}})\in\mathbb{U}\times\mathbb{R}^{l}$ are the augmented inputs.
Theorem 3.7 (DSMs are CBFs).
If $\Delta:\mathcal{D}_{\boldsymbol{\mathbf{v}}}\times\mathbb{V}\to\mathbb{R}$ is a DSM for $\pi$, then it is a CBF with respect to the augmented system (22).
Proof 3.8.
Given $\alpha\in\mathscr{K}_{\infty}$, consider $(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\mathcal{D}_{\boldsymbol{\mathbf{v}}}\times\mathbb{V}$ such that $\Delta(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\geq 0$. By (19b), $\boldsymbol{\mathbf{x}}\in\mathcal{X}_{\boldsymbol{\mathbf{v}}}$, which implies that $\pi(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\mathcal{U}$. Consider the augmented input $\big{(}\pi(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}}),0\big{)}\in\mathcal{U}\times\mathbb{R}^{l}$ and note that, by (19c),
$\dot{\Delta}\big{(}\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}},\pi(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}}),0\big{)}$ $=\frac{\partial\Delta}{\partial\boldsymbol{\mathbf{x}}}\big{(}f(\boldsymbol{\mathbf{x}})+g(\boldsymbol{\mathbf{x}})\pi(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\big{)}+\frac{\partial\Delta}{\partial\boldsymbol{\mathbf{v}}}0$
$=\frac{\partial\Delta}{\partial\boldsymbol{\mathbf{x}}}f_{\pi}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})$
$\geq 0$
$\geq-\alpha\big{(}\Delta(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\big{)}.$
It then follows that, $\forall(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\mathcal{D}_{\boldsymbol{\mathbf{v}}}\times\mathbb{V}$ such that $\Delta(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\geq 0$,
$\sup_{(\boldsymbol{\mathbf{u}},\boldsymbol{\mathbf{w}})\in\mathcal{U}\times\mathbb{R}^{l}}\left[\dot{\Delta}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}},\boldsymbol{\mathbf{u}},\boldsymbol{\mathbf{w}})\right]\geq-\alpha\big{(}\Delta(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\big{)}.$ (23)
Therefore, $\Delta(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})$ is a CBF.
For DSMs, the sets $\mathcal{C}\subset\mathbb{X}\times\mathbb{V}$ and $\mathcal{K}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\subset\mathbb{U}\times\mathbb{R}^{l}$ are
$\mathcal{C}$ =$=\{(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\mathcal{D}_{\boldsymbol{\mathbf{v}}}\!\times\!\mathbb{V}\mid\Delta(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\geq 0\}\subset\mathcal{X}\times\mathcal{V},$ (24)
$\mathcal{K}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})$ =$=\{(\boldsymbol{\mathbf{u}},\boldsymbol{\mathbf{w}})\in\mathcal{U}\times\mathbb{R}^{l}\mid\dot{\Delta}+\alpha\big{(}\Delta(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\big{)}\geq 0\}.$ (25)
A significant concern in the CBF literature arises when the term $L_{g}h(\boldsymbol{\mathbf{x}})=0$ for some $\boldsymbol{\mathbf{x}}\in\mathcal{C}$. This means $h$ lacks uniform relative degree 1 over $\mathcal{C}$, and implies a loss of controllability that can be catastrophic when $L_{f}h(\boldsymbol{\mathbf{x}})<-\alpha\big{(}h(\boldsymbol{\mathbf{x}})\big{)}$. Although many relevant works address this issue (e.g. [7, 8]), the following lemma shows that $L_{g}h(\boldsymbol{\mathbf{x}})=0$ is never a concern for DSM-based CBFs.
Lemma 3.9.
Let $\Delta$ be a DSM for $\pi$ and let $(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\mathcal{C}$ be such that $\begin{bmatrix}\frac{\partial\Delta}{\partial\boldsymbol{\mathbf{x}}}g(\boldsymbol{\mathbf{x}})&\frac{\partial\Delta}{\partial\boldsymbol{\mathbf{v}}}\end{bmatrix}=0$. Then, $\mathcal{K}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})=\mathcal{U}\times\mathbb{R}^{l}$.
Proof 3.10.
By construction, $\big{(}\pi(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}}),0\big{)}\in\mathcal{K}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})$. Thus,
$\dot{\Delta}\big{(}\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}},\pi(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}}),0\big{)}$ $=\frac{\partial\Delta}{\partial\boldsymbol{\mathbf{x}}}f(\boldsymbol{\mathbf{x}})+\begin{bmatrix}\frac{\partial\Delta}{\partial\boldsymbol{\mathbf{x}}}g(\boldsymbol{\mathbf{x}})&\frac{\partial\Delta}{\partial\boldsymbol{\mathbf{v}}}\end{bmatrix}\begin{bmatrix}\pi(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\\ 0\end{bmatrix}$
$=\frac{\partial\Delta}{\partial\boldsymbol{\mathbf{x}}}f(\boldsymbol{\mathbf{x}})\geq-\alpha\big{(}\Delta(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\big{)}.$
Therefore, any admissible augmented input $(\boldsymbol{\mathbf{u}},\boldsymbol{\mathbf{w}})\in\mathcal{U}\times\mathbb{R}^{l}$ trivially satisfies the CBF condition.
Recalling Theorem 3.7, it is possible to reformulate the CBF policy $\beta$ for the augmented system $(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})$. To do so, let $\boldsymbol{\mathbf{r}}\in\mathbb{V}$ be a desired reference, let $\kappa:\mathbb{X}\times\mathbb{V}\to\mathbb{U}$ be a nominal control policy with desirable closed-loop performance, and let $\rho:\mathbb{V}\times\mathbb{V}\to\mathbb{R}^{l}$ be a navigation field. The following theorem formulates the CBF-based policy $\beta$ for DSMs and the augmented system (22).
Theorem 3.11 (DSM-CBF policy).
Let $\Delta$ be a DSM and let $\mathcal{C}$, given in (24), be compact. Assume also that 00 is a regular value of $\Delta$ and let $\eta>0$ be a constant scalar. If the optimization-based policy $\beta:\mathcal{C}\to\mathcal{U}\times\mathbb{R}^{l}$
$\beta(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})=\!\underset{(\boldsymbol{\mathbf{u}},\boldsymbol{\mathbf{w}})\in\mathcal{K}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})}{\operatorname{argmin}}\|\boldsymbol{\mathbf{u}}\!-\!\kappa(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{r}})\|^{2}\!+\eta\|\boldsymbol{\mathbf{w}}\!-\!\rho(\boldsymbol{\mathbf{v}},\boldsymbol{\mathbf{r}})\|^{2},$ (26)
is locally Lipschitz continuous, then the trajectory of the augmented system (22) under policy $(\boldsymbol{\mathbf{u}},\boldsymbol{\mathbf{w}})=\beta(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})$ exists, is unique, and satisfies $\boldsymbol{\mathbf{x}}(t)\in\mathcal{X}$ and $\boldsymbol{\mathbf{v}}(t)\in\mathcal{V}$ for all $t\geq 0$, given $\big{(}\boldsymbol{\mathbf{x}}(0),\boldsymbol{\mathbf{v}}(0)\big{)}\in\mathcal{C}$.
Proof 3.12.
By Theorem 3.7, $\Delta$ is a CBF and, since the set $\mathcal{K}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\neq\emptyset$, compactness on $\mathcal{U}$ is not needed to apply Lemma 1. So, the map $\beta$ is well-defined. Local Lipschitzness of $\beta$ guarantees existence and uniqueness of the solution to the closed-loop system. Safety follows immediately because regularity of 00 in $\Delta$ and the condition $\dot{\Delta}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}},\boldsymbol{\mathbf{u}},\boldsymbol{\mathbf{w}})\geq-\alpha\big{(}\Delta(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\big{)}$, ensure invariance of $\mathcal{C}\subset\mathcal{X}\times\mathcal{V}$ for the dynamics (22) with input $(\boldsymbol{\mathbf{u}},\boldsymbol{\mathbf{w}})=\beta(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})$.
Remark 3.13.
Although the set $\mathcal{K}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})$ is generally not convex, if the input constraint set $\mathcal{U}$ is polyhedral, the optimization problem (26) becomes a quadratic program (QP) for all $(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\mathcal{C}$, which can be solved efficiently.
Remark 3.14.
The positive constant $\eta>0$ makes the objective function of the augmented CBF-program (26) strictly convex. Moreover, we observed in practice that as $\eta\to\infty$, the system behavior approaches that of the ERG, which tends to have worse performance than can be achieved with $\eta\to 0$.
Remark 3.15.
If $\mathcal{V}$ is convex, a suitable navigation field is $\rho(\boldsymbol{\mathbf{v}},\boldsymbol{\mathbf{r}})=\boldsymbol{\mathbf{r}}-\boldsymbol{\mathbf{v}}$. Otherwise, please refer to works like [32].
Remark 3.16.
The proposed approach can be interpreted as a closed-form variant of the backup CBF [20]. The approach assigns a safe control input $\boldsymbol{\mathbf{u}}$ while simultaneously generating a safe reference $\boldsymbol{\mathbf{v}}$ for the backup controller $\pi$.
Remark 3.17.
Note that the nominal controller $\kappa$ and the prestabilizing controller $\pi$ can be different. In practice, we observed that mild prestabilizing controllers $\pi$ lead to larger feasible regions on the CBF-based program, allowing the safe input to be closer to the nominal target $\kappa$.
We provide some conditions that guarantee the CBF policy $\beta$ for DSM-based CBFs is locally Lipschitz continuous. Let us assume the input constraint set is polyhedral $\mathcal{U}=\{\boldsymbol{\mathbf{u}}\in\mathbb{U}\mid M\boldsymbol{\mathbf{u}}\leq\boldsymbol{\mathbf{b}}\}$. We begin by defining the following linear program (LP), which finds the “width” of the feasible set [33]
$\omega^{*}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})=$ $\max_{(\boldsymbol{\mathbf{u}},\boldsymbol{\mathbf{w}},\omega)\in\mathbb{U}\times\mathbb{R}^{l}\times\mathbb{R}}\quad\omega$ (27a)
s.t. $\begin{bmatrix}A(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\end{bmatrix}\begin{bmatrix}\boldsymbol{\mathbf{u}}\\ \boldsymbol{\mathbf{w}}\end{bmatrix}+\boldsymbol{\mathbf{1}}\omega\leq b(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}}),$ (27b)
where $\boldsymbol{\mathbf{1}}$ is a column vector of $1$’s,
$A(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})=\begin{bmatrix}-\frac{\partial\Delta}{\partial\boldsymbol{\mathbf{x}}}g(\boldsymbol{\mathbf{x}})&-\frac{\partial\Delta}{\partial\boldsymbol{\mathbf{v}}}\\ M&0\end{bmatrix},$
and
$b(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})=\begin{bmatrix}\alpha\big{(}\Delta(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\big{)}+\frac{\partial\Delta}{\partial\boldsymbol{\mathbf{x}}}f(\boldsymbol{\mathbf{x}})\\ \boldsymbol{\mathbf{b}}\end{bmatrix}.$
Based on [33], we can formulate a reduced set of sufficient conditions that guarantee $\beta$ is locally Lipschitz continuous.
Theorem 3.18.
[33][Theorem 1] Consider $\beta(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})$ defined as in (26) and suppose the following conditions hold at a point $(\boldsymbol{\mathbf{x}}_{0},\boldsymbol{\mathbf{v}}_{0})\in\mathcal{C}$:
  1. 1.
    $\omega^{*}(\boldsymbol{\mathbf{x}}_{0},\boldsymbol{\mathbf{v}}_{0})>0$,
  2. 2.
    $\frac{\partial\Delta}{\partial\boldsymbol{\mathbf{x}}}$ and $\frac{\partial\Delta}{\partial\boldsymbol{\mathbf{v}}}$ are locally Lipschitz continuous at $(\boldsymbol{\mathbf{x}}_{0},\boldsymbol{\mathbf{v}}_{0})$.
Then, the feedback $\beta(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})$ is unique and locally Lipschitz continuous at $(\boldsymbol{\mathbf{x}}_{0},\boldsymbol{\mathbf{v}}_{0})$.
The following lemma guarantees that $\omega^{*}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})>0$ wherever the feasible set has a nonempty interior.
Lemma 3.19.
If $\operatorname{Int}\left(\mathcal{K}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\right)\neq\emptyset$, then $\omega^{*}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})>0$.
Proof 3.20.
Note that $\mathcal{K}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})=\{(\boldsymbol{\mathbf{u}},\boldsymbol{\mathbf{w}})\in\mathbb{U}\times\mathbb{R}^{l}:A(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})[\boldsymbol{\mathbf{u}};~{}\boldsymbol{\mathbf{w}}]\leq b(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\}$. Let $(\boldsymbol{\mathbf{u}}_{0},\boldsymbol{\mathbf{w}}_{0})\in\operatorname{Int}\left(\mathcal{K}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\right)$, that is, $A(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})[\boldsymbol{\mathbf{u}}_{0};\boldsymbol{\mathbf{w}}_{0}]<b(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})$. It follows that $\exists\omega_{0}>0$ such that $A(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})[\boldsymbol{\mathbf{u}}_{0};\boldsymbol{\mathbf{w}}_{0}]+\boldsymbol{\mathbf{1}}\omega_{0}\leq b(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})$. The point $(\boldsymbol{\mathbf{u}}_{0},\boldsymbol{\mathbf{w}}_{0},\omega_{0})$ is feasible in the LP defining $\omega^{*}$. Thus, $\omega^{*}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\geq\omega_{0}>0$.
4 Control-sharing Property
Often, the sets $\mathcal{X}$ and $\mathcal{U}$ are the intersection of multiple state and input constraints. In these cases, it is easier to design a separate CBF for each constraint and then enforce all the conditions in one CBF-based program. Unfortunately, there is typically no guarantee that feasibility is retained when all constraints are superimposed. This section shows that DSMs designed for the same prestabilized system have the control-sharing property [9], which makes their CBF conditions compatible.
Consider again the control-affine system (1). Given $q\in\mathbb{N}$ distinct CBFs $h_{i}:\mathcal{D}_{i}\to\mathbb{R}$ and their associated control invariant sets $\mathcal{C}_{i}=\{\boldsymbol{\mathbf{x}}\in\mathcal{D}_{i}\mid h_{i}(\boldsymbol{\mathbf{x}})\geq 0\}$, define the $i$-th set of safe inputs as
$\mathcal{K}_{i}(\boldsymbol{\mathbf{x}})=\left\{\boldsymbol{\mathbf{u}}\in\mathcal{U}\mid L_{f}h_{i}(\boldsymbol{\mathbf{x}})+L_{g}h_{i}(\boldsymbol{\mathbf{x}})\boldsymbol{\mathbf{u}}+\alpha_{i}\big{(}h_{i}(\boldsymbol{\mathbf{x}})\big{)}\geq 0\right\},$
where each $\alpha_{i}$ is a class $\mathscr{K}_{\infty}$ function associated with $h_{i}$. If the conditions of Lemma 1 hold for each $h_{i}$, we have that
$\forall i\in\{1,\ldots,q\},\quad\boldsymbol{\mathbf{x}}\in\mathcal{C}_{i}\implies\mathcal{K}_{i}(\boldsymbol{\mathbf{x}})\neq\emptyset.$ (28)
However, there is generally no guarantee that the intersection $\bigcap_{i=1}^{q}\mathcal{K}_{i}(\boldsymbol{\mathbf{x}})$ is nonempty.
Definition 4.21.
[9, Definition 3] Consider $q\in\mathbb{N}$ CBFs $h_{i}:\mathcal{D}_{i}\to\mathbb{R}$. These CBFs are said to have the control-sharing property if
$\forall\boldsymbol{\mathbf{x}}\in\bigcap_{i=1}^{q}\mathcal{C}_{i},\quad\bigcap_{i=1}^{q}\mathcal{K}_{i}(\boldsymbol{\mathbf{x}})\neq\emptyset.$ (29)
If the CBFs $h_{i}$ have the control-sharing property, then the optimization-based control policy $\beta$ defined in (12) with
$\mathcal{C}=\bigcap_{i=1}^{q}\mathcal{C}_{i},\quad\mathcal{K}(\boldsymbol{\mathbf{x}})=\bigcap_{i=1}^{q}\mathcal{K}_{i}(\boldsymbol{\mathbf{x}}),$ (30)
is feasible for all $\boldsymbol{\mathbf{x}}\in\mathcal{C}$. If it is also locally Lipschitz continuous, then $\boldsymbol{\mathbf{x}}(t)\in\mathcal{C}$ for all $t\geq 0$.
Now, suppose the state $\mathcal{X}$ and input $\mathcal{U}$ constraint sets can be decomposed into $q\in\mathbb{N}$ closed constraint sets $\mathcal{X}_{i}\subset\mathbb{X}$ and $\mathcal{U}_{i}\subset\mathbb{U}$ such that $\mathcal{X}\times\mathcal{U}\subset\mathcal{X}_{i}\times\mathcal{U}_{i},~{}\forall i\in\{1,\ldots,q\},$ and $\bigcap_{i=1}^{q}\mathcal{X}_{i}\times\mathcal{U}_{i}=\mathcal{X}\times\mathcal{U}$. The $i$-th steady-state admissible set is $\mathcal{V}_{i}=\{\boldsymbol{\mathbf{v}}\in\mathbb{V}\mid\bar{x}(\boldsymbol{\mathbf{v}})\in\mathcal{X}_{i},~{}\bar{u}(\boldsymbol{\mathbf{r}})\in\mathcal{U}_{i}\}$ and note that $\bigcap_{i=1}^{q}\mathcal{V}_{i}=\mathcal{V}$. Let $\pi:\mathbb{X}\times\mathbb{V}\to\mathbb{U}$ be a prestabilizing controller as described above and assume that $\Delta_{i}:\mathcal{D}_{\boldsymbol{\mathbf{v}}}\times\mathbb{V}\to\mathbb{R}$ are dynamic safety margins for $\pi$ with respect to the constraint sets $\mathcal{X}_{i}$, $\mathcal{U}_{i}$. The following theorem shows that these DSMs have the control-sharing property.
Theorem 4.22.
Any number of DSMs $\Delta_{i}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})$ for the same prestabilizing policy $\pi$ have the control-sharing property.
Proof 4.23.
Let $q\in\mathbb{N}$ and consider $q$ DSMs, $\Delta_{i}$, for $\pi$. If $\mathcal{C}=\bigcap_{i=1}^{q}\mathcal{C}_{i}=\emptyset$, Definition 4.21 holds trivially. Otherwise, let $(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\mathcal{C}$ and $i\in\{1,\ldots,q\}$ be given. Since $\Delta_{i}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\geq 0$, it follows by (19b) that $\pi(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\mathcal{U}_{i}$. For any $\alpha\in\mathscr{K}_{\infty}$, it can be shown that the augmented input $\big{(}\pi(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}}),0\big{)}\in\mathcal{K}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})$. Since $i$ was arbitrary, it follows that $\mathcal{K}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})=\bigcap_{i=1}^{q}\mathcal{K}_{i}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\neq\emptyset$ and we conclude that the DSMs $\Delta_{i}$ have the control-sharing property.
The next section provides a specific method for constructing DSMs using Lyapunov functions.
5 Lyapunov-based Dynamic Safety Margins
This section shows that the Lyapunov-based DSMs in [22, 28, 29, 30] can be used within the context of this paper. Let a continuously differentiable function $V:\mathbb{X}\times\mathbb{V}\to\mathbb{R}$ be a reference-dependent Lyapunov function for the prestabilized system $f_{\pi}$. That is, for all $\boldsymbol{\mathbf{v}}\in\mathbb{V}$,
$V\big{(}\bar{x}(\boldsymbol{\mathbf{v}}),\boldsymbol{\mathbf{v}}\big{)}$ $=0,$ (31a)
$V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})$ $>0,\quad\forall\boldsymbol{\mathbf{x}}\in\mathcal{D}_{\boldsymbol{\mathbf{v}}}\setminus\{\bar{x}(\boldsymbol{\mathbf{v}})\},$ (31b)
$\frac{\partial V}{\partial\boldsymbol{\mathbf{x}}}f_{\pi}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})$ $\leq 0,\quad\forall\boldsymbol{\mathbf{x}}\in\mathcal{D}_{\boldsymbol{\mathbf{v}}}.$ (31c)
Additionally, we require $V$ to be strictly monotonically increasing over the sets $\mathcal{D}_{\boldsymbol{\mathbf{v}}}$. Formally, we have the following assumption.
Assumption 4
Given any $\boldsymbol{\mathbf{v}}\in\mathbb{V}$, $\boldsymbol{\mathbf{x}}\in\mathcal{D}_{\boldsymbol{\mathbf{v}}}\setminus\{\bar{x}(\boldsymbol{\mathbf{v}})\}$ is a regular point of $V(\cdot,\boldsymbol{\mathbf{v}}):\mathbb{X}\to\mathbb{R}$.
Recalling the set $\mathcal{X}_{\boldsymbol{\mathbf{v}}}$ defined in (18), let $\mathcal{X}^{c}_{\boldsymbol{\mathbf{v}}}=\mathbb{X}\setminus\mathcal{X}_{\boldsymbol{\mathbf{v}}}$ be its complement, which is open in $\mathbb{X}$. We define $\Gamma^{*}:\mathbb{V}\to\mathbb{R}$ as
$\Gamma^{*}(\boldsymbol{\mathbf{v}})\triangleq\inf_{\boldsymbol{\mathbf{x}}\in\>\mathcal{X}^{c}_{\boldsymbol{\mathbf{v}}}\cap\mathcal{D}_{\boldsymbol{\mathbf{v}}}}V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})-\inf_{\boldsymbol{\mathbf{x}}\in\mathcal{X}_{\boldsymbol{\mathbf{v}}}\cap\mathcal{D}_{\boldsymbol{\mathbf{v}}}}V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}}).$ (32)
When positive, $\Gamma^{*}(\boldsymbol{\mathbf{v}})$ represents the largest constraint-admissible level set of the Lyapunov function $V$ at reference $\boldsymbol{\mathbf{v}}$. Our definition for $\Gamma^{*}$ is an extension of the threshold value in [22] to the entire reference space $\mathbb{V}$. With this, $\Gamma^{*}$ has the following property.
Lemma 5.24.
Let $V:\mathbb{X}\times\mathbb{V}\to\mathbb{R}$ be a reference-dependent Lyapunov function that satisfies Assumption 4. We have that
$\Gamma^{*}(\boldsymbol{\mathbf{v}})\geq 0\quad\iff\quad\boldsymbol{\mathbf{v}}\in\mathcal{V}.$ (33)
Proof 5.25.
Let $\boldsymbol{\mathbf{v}}\in\mathcal{V}$ be given. Then, $\bar{x}(\boldsymbol{\mathbf{v}})\in\mathcal{X}_{\boldsymbol{\mathbf{v}}}\cap\mathcal{D}_{\boldsymbol{\mathbf{v}}}$, which implies that $\inf_{\boldsymbol{\mathbf{x}}\in\mathcal{X}_{\boldsymbol{\mathbf{v}}}\cap\mathcal{D}_{\boldsymbol{\mathbf{v}}}}V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})=0$. Recalling that $\forall\boldsymbol{\mathbf{x}}\in\mathcal{D}_{\boldsymbol{\mathbf{v}}}$, $V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\geq 0$, we conclude that $\Gamma^{*}(\boldsymbol{\mathbf{v}})\geq 0$.
For the other direction, we show the contrapositive statement. Let $\boldsymbol{\mathbf{v}}\in\mathbb{V}\setminus\mathcal{V}$ be given. Then, $\bar{x}(\boldsymbol{\mathbf{v}})\notin\mathcal{X}_{\boldsymbol{\mathbf{v}}}$, and $\bar{x}(\boldsymbol{\mathbf{v}})\in\mathcal{X}_{\boldsymbol{\mathbf{v}}}^{c}\cap\mathcal{D}_{\boldsymbol{\mathbf{v}}}$, which implies that $\inf_{\boldsymbol{\mathbf{x}}\in\mathcal{X}_{\boldsymbol{\mathbf{v}}}^{c}\cap\mathcal{D}_{\boldsymbol{\mathbf{v}}}}V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})=0$. Since $\mathcal{X}_{\boldsymbol{\mathbf{v}}}^{c}\cap\mathcal{D}_{\boldsymbol{\mathbf{v}}}$ is open in $\mathbb{X}$, there exists $\delta>0$ such that $\mathcal{B}\big{(}\bar{x}(\boldsymbol{\mathbf{v}}),\delta\big{)}=\{\boldsymbol{\mathbf{x}}\in\mathbb{X}:\|\boldsymbol{\mathbf{x}}-\bar{x}(\boldsymbol{\mathbf{v}})\|<\delta\}\subset\mathcal{X}_{\boldsymbol{\mathbf{v}}}^{c}\cap\mathcal{D}_{\boldsymbol{\mathbf{v}}}$. Furthermore, by compactness of $\partial\mathcal{B}\big{(}\bar{x}(\boldsymbol{\mathbf{v}}),\delta\big{)}$, there exists $\boldsymbol{\mathbf{x}}_{0}\in\partial\mathcal{B}\big{(}\bar{x}(\boldsymbol{\mathbf{v}}),\delta\big{)}$, with $\boldsymbol{\mathbf{x}}_{0}\neq\bar{x}(\boldsymbol{\mathbf{v}})$ because $\bar{x}(\boldsymbol{\mathbf{v}})\notin\partial\mathcal{B}\big{(}\bar{x}(\boldsymbol{\mathbf{v}}),\delta\big{)}$, such that
$\inf_{\boldsymbol{\mathbf{x}}\in\partial\mathcal{B}(\bar{x}(\boldsymbol{\mathbf{v}}),\delta)}V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})=V(\boldsymbol{\mathbf{x}}_{0},\boldsymbol{\mathbf{v}})>0.$ (34)
By Assumption 4, $V$ is monotonically increasing over $\mathcal{D}_{\boldsymbol{\mathbf{v}}}$, and its level sets contained in $\mathcal{D}_{\boldsymbol{\mathbf{v}}}$ are connected. It follows that $\mathcal{L}\triangleq\{\boldsymbol{\mathbf{x}}\in\mathcal{D}_{\boldsymbol{\mathbf{v}}}:V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\leq V(\boldsymbol{\mathbf{x}}_{0},\boldsymbol{\mathbf{v}})\}\subset\mathcal{B}\big{(}\bar{x}(\boldsymbol{\mathbf{v}}),\delta\big{)}$. Now, let $\boldsymbol{\mathbf{x}}\in\mathcal{X}_{\boldsymbol{\mathbf{v}}}\cap\mathcal{D}_{\boldsymbol{\mathbf{v}}}$ be given and note that $\boldsymbol{\mathbf{x}}\notin\mathcal{B}\big{(}\bar{x}(\boldsymbol{\mathbf{v}}),\delta\big{)}$, implying that $\boldsymbol{\mathbf{x}}\notin\mathcal{L}$. So, we must have that $V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})>V(\boldsymbol{\mathbf{x}}_{0},\boldsymbol{\mathbf{v}})$. As a result,
$\inf_{\boldsymbol{\mathbf{x}}\in\mathcal{X}_{\boldsymbol{\mathbf{v}}}\cap\mathcal{D}_{\boldsymbol{\mathbf{v}}}}V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\geq V(\boldsymbol{\mathbf{x}}_{0},\boldsymbol{\mathbf{v}})>0,$ (35)
and we conclude that $\Gamma^{*}(\boldsymbol{\mathbf{v}})\leq-V(\boldsymbol{\mathbf{x}}_{0},\boldsymbol{\mathbf{v}})<0$.
Let us define the stability threshold value
$\overline{\Gamma}(\boldsymbol{\mathbf{v}})\triangleq\inf_{\boldsymbol{\mathbf{x}}\in\partial\mathcal{D}_{\boldsymbol{\mathbf{v}}}}V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}}).$ (36)
We can now construct DSMs using Lyapunov functions.
Theorem 5.26 (Lyapunov-based DSMs).
Let $V:\mathbb{X}\times\mathbb{V}\to\mathbb{R}$ be a reference-dependent Lyapunov function for the prestabilized system $f_{\pi}$ satisfying Assumption 4. Let $\Gamma:\mathbb{V}\to\mathbb{R}$ be a continuously differentiable function such that $\forall\boldsymbol{\mathbf{v}}\in\mathcal{V}$, $0\leq\Gamma(\boldsymbol{\mathbf{v}})\leq\min\left(\Gamma^{*}(\boldsymbol{\mathbf{v}}),\overline{\Gamma}(\boldsymbol{\mathbf{v}})\right)$ and $\forall\boldsymbol{\mathbf{v}}\in\mathbb{V}\setminus\mathcal{V}$, $\Gamma(\boldsymbol{\mathbf{v}})<0$. Then, $\Delta:\mathcal{D}_{\boldsymbol{\mathbf{v}}}\times\mathbb{V}\to\mathbb{R}$, defined as
$\Delta(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})=\Gamma(\boldsymbol{\mathbf{v}})-V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}}),$ (37)
is a dynamic safety margin for $\pi$.
Proof 5.27.
To prove property (19a) in Definition 2.4, let $\boldsymbol{\mathbf{v}}\in\mathcal{V}$. Then, by Lemma 5.24, $\Delta\big{(}\bar{x}(\boldsymbol{\mathbf{v}}),\boldsymbol{\mathbf{v}}\big{)}=\Gamma(\boldsymbol{\mathbf{v}})-V\big{(}\bar{x}(\boldsymbol{\mathbf{v}}),\boldsymbol{\mathbf{v}}\big{)}=\Gamma(\boldsymbol{\mathbf{v}})\geq 0$.
To prove property (19b), let $(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\mathcal{D}_{\boldsymbol{\mathbf{v}}}\times\mathbb{V}$ be such that $\Delta(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\geq 0$. For a contradiction, assume $\boldsymbol{\mathbf{x}}\notin\mathcal{X}_{\boldsymbol{\mathbf{v}}}$. Then
$\Gamma^{*}(\boldsymbol{\mathbf{v}})\leq V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})-\inf_{\boldsymbol{\mathbf{x}}\in\mathcal{X}_{\boldsymbol{\mathbf{v}}}\cap\mathcal{D}_{\boldsymbol{\mathbf{v}}}}V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\leq V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}}).$
Additionally, $\Delta(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\geq 0$ implies that $\Gamma^{*}(\boldsymbol{\mathbf{v}})\geq\Gamma(\boldsymbol{\mathbf{v}})\geq V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\geq 0$, which further implies by Lemma 5.24 that $\boldsymbol{\mathbf{v}}\in\mathcal{V}$. Thus,
$\Gamma^{*}(\boldsymbol{\mathbf{v}})\leq V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\leq\Gamma(\boldsymbol{\mathbf{v}})\leq\Gamma^{*}(\boldsymbol{\mathbf{v}}).$ (38)
So, we must have $\Gamma(\boldsymbol{\mathbf{v}})=V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})=\Gamma^{*}(\boldsymbol{\mathbf{v}})$. If $V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})=0$, then $\boldsymbol{\mathbf{x}}=\bar{x}(\boldsymbol{\mathbf{v}})\in\mathcal{X}_{\boldsymbol{\mathbf{v}}}$ because $\boldsymbol{\mathbf{v}}\in\mathcal{V}$, and we have a contradiction. Otherwise, if $\Gamma^{*}(\boldsymbol{\mathbf{v}})=V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})>0$, recall that $\mathcal{X}^{c}_{\boldsymbol{\mathbf{v}}}\cap\mathcal{D}_{\boldsymbol{\mathbf{v}}}$ is open in $\mathbb{X}$. Therefore, there exists an open set $\mathcal{N}\subset\mathcal{X}^{c}_{\boldsymbol{\mathbf{v}}}\cap\mathcal{D}_{\boldsymbol{\mathbf{v}}}$ with $\boldsymbol{\mathbf{x}}\in\mathcal{N}$. Furthermore, since $\boldsymbol{\mathbf{x}}\in\mathcal{D}_{\boldsymbol{\mathbf{v}}}\setminus\{\bar{x}(\boldsymbol{\mathbf{v}})\}$ is a regular point of $V(\cdot,\boldsymbol{\mathbf{v}})$ by Assumption 4, we have that $\frac{\partial V}{\partial\boldsymbol{\mathbf{x}}}\neq 0$. So, there exists a sufficiently small scalar $\lambda>0$ such that $\tilde{\boldsymbol{\mathbf{x}}}\triangleq\boldsymbol{\mathbf{x}}-\lambda\frac{\partial V}{\partial\boldsymbol{\mathbf{x}}}^{\top}\in\mathcal{N}\subset\mathcal{X}^{c}_{\boldsymbol{\mathbf{v}}}\cap\mathcal{D}_{\boldsymbol{\mathbf{v}}}$. Furthermore, this point achieves lower value in $V$. That is, $V(\tilde{\boldsymbol{\mathbf{x}}},\boldsymbol{\mathbf{v}})<V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})$. Since $\boldsymbol{\mathbf{v}}\in\mathcal{V}$, we have that $\bar{x}(\boldsymbol{\mathbf{v}})\in\mathcal{X}_{\boldsymbol{\mathbf{v}}}\cap\mathcal{D}_{\boldsymbol{\mathbf{v}}}$. Thus, $\Gamma^{*}(\boldsymbol{\mathbf{v}})=\inf_{\boldsymbol{\mathbf{x}}\in\mathcal{X}_{\boldsymbol{\mathbf{v}}}^{c}\cap\mathcal{D}_{\boldsymbol{\mathbf{v}}}}V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})$. So, $\Gamma^{*}(\boldsymbol{\mathbf{v}})=V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})$ is a contradiction because
$\Gamma^{*}(\boldsymbol{\mathbf{v}})=\inf_{\boldsymbol{\mathbf{x}}\in\mathcal{X}_{\boldsymbol{\mathbf{v}}}\cap\mathcal{D}_{\boldsymbol{\mathbf{v}}}}V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\leq V(\tilde{\boldsymbol{\mathbf{x}}},\boldsymbol{\mathbf{v}})<V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}}),$
and we must have that $\boldsymbol{\mathbf{x}}\in\mathcal{X}_{\boldsymbol{\mathbf{v}}}$.
To prove property (19c), let $(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\mathcal{D}_{\boldsymbol{\mathbf{v}}}\times\mathbb{V}$ be such that $\Delta(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\geq 0$. By (31c),
$\frac{\partial\Delta}{\partial\boldsymbol{\mathbf{x}}}f_{\pi}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})=-\frac{\partial V}{\partial\boldsymbol{\mathbf{x}}}f_{\pi}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\geq 0.$ (39)
To prove the final property (19d), let $\boldsymbol{\mathbf{v}}\in\mathbb{V}$ and $\epsilon>0$ be given. If $\boldsymbol{\mathbf{v}}\notin\mathcal{V}$, then $\Gamma(\boldsymbol{\mathbf{v}})<0$ and $V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\geq 0$ for all $\boldsymbol{\mathbf{x}}\in\mathcal{D}_{\boldsymbol{\mathbf{v}}}$. Thus, $\Delta(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})<0$ for all $\boldsymbol{\mathbf{x}}\in\mathcal{D}_{\boldsymbol{\mathbf{v}}}$ and we have that $\mathcal{C}_{\boldsymbol{\mathbf{v}}}^{\epsilon}=\emptyset$ which is trivially closed in $\mathbb{X}$. Otherwise, if $\boldsymbol{\mathbf{v}}\in\mathcal{V}$, let $\{\boldsymbol{\mathbf{x}}_{n}\}\subset\mathcal{C}_{\boldsymbol{\mathbf{v}}}^{\epsilon}$ be a sequence that converges to $\tilde{\boldsymbol{\mathbf{x}}}\in\mathbb{X}$. Define the map $\zeta:\boldsymbol{\mathbf{x}}\mapsto\Gamma(\boldsymbol{\mathbf{v}})-V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})$. By continuity of $\zeta$ on $\mathbb{X}$, we have that the set $\mathcal{S}=\{\boldsymbol{\mathbf{x}}\in\mathbb{X}:\zeta(\boldsymbol{\mathbf{x}})\geq\epsilon\}$ is closed in $\mathbb{X}$. Since $\{\boldsymbol{\mathbf{x}}_{n}\}\subset\mathcal{C}_{\boldsymbol{\mathbf{v}}}^{\epsilon}\subset\mathcal{S}$, we have that $\tilde{\boldsymbol{\mathbf{x}}}\in\mathcal{S}$. It remains to show that $\tilde{\boldsymbol{\mathbf{x}}}\in\mathcal{D}_{\boldsymbol{\mathbf{v}}}$. Begin by noting that $\zeta(\tilde{\boldsymbol{\mathbf{x}}})=\Gamma(\boldsymbol{\mathbf{v}})-V(\tilde{\boldsymbol{\mathbf{x}}},\boldsymbol{\mathbf{v}})\geq\epsilon$ because $\tilde{\boldsymbol{\mathbf{x}}}\in\mathcal{S}$. Thus,
$\min\big{(}\Gamma^{*}(\boldsymbol{\mathbf{v}}),\overline{\Gamma}(\boldsymbol{\mathbf{v}})\big{)}\geq\Gamma(\boldsymbol{\mathbf{v}})\geq V(\tilde{\boldsymbol{\mathbf{x}}},\boldsymbol{\mathbf{v}})+\epsilon.$
For a contradiction, assume $\tilde{\boldsymbol{\mathbf{x}}}\in\mathcal{D}^{c}_{\boldsymbol{\mathbf{v}}}=\mathbb{X}\setminus\mathcal{D}_{\boldsymbol{\mathbf{v}}}$. If $\tilde{\boldsymbol{\mathbf{x}}}\in\partial\mathcal{D}_{\boldsymbol{\mathbf{v}}}\subset\mathcal{D}^{c}_{\boldsymbol{\mathbf{v}}}$, then we have that
$\overline{\Gamma}(\boldsymbol{\mathbf{v}})=\inf_{\boldsymbol{\mathbf{x}}\in\partial\mathcal{D}_{\boldsymbol{\mathbf{v}}}}V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\geq V(\tilde{\boldsymbol{\mathbf{x}}},\boldsymbol{\mathbf{v}})+\epsilon>V(\tilde{\boldsymbol{\mathbf{x}}},\boldsymbol{\mathbf{v}}),$
which is a contradiction. On the other hand, if $\tilde{\boldsymbol{\mathbf{x}}}\in\operatorname{Int}\left(\mathcal{D}^{c}_{\boldsymbol{\mathbf{v}}}\right)\subset\mathcal{D}^{c}_{\boldsymbol{\mathbf{v}}}$, there exists an open set $\mathcal{N}\subset\operatorname{Int}\left(\mathcal{D}^{c}_{\boldsymbol{\mathbf{v}}}\right)$ with $\tilde{\boldsymbol{\mathbf{x}}}\in\mathcal{N}$. By convergence of $\{\boldsymbol{\mathbf{x}}_{n}\}$ to $\tilde{\boldsymbol{\mathbf{x}}}$, there exists $N\in\mathbb{N}$ such that $\boldsymbol{\mathbf{x}}_{N}\in\mathcal{N}\subset\operatorname{Int}\left(\mathcal{D}^{c}_{\boldsymbol{\mathbf{v}}}\right)\subset\mathcal{D}^{c}_{\boldsymbol{\mathbf{v}}}$. This is a contradiction because $\boldsymbol{\mathbf{x}}_{N}\in\mathcal{D}_{\boldsymbol{\mathbf{v}}}$. So, we must have that $\tilde{\boldsymbol{\mathbf{x}}}\in\mathcal{D}_{\boldsymbol{\mathbf{v}}}\cap\mathcal{S}=\mathcal{C}_{\boldsymbol{\mathbf{v}}}^{\epsilon}$. It follows that $\mathcal{C}_{\boldsymbol{\mathbf{v}}}^{\epsilon}$ is closed in $\mathbb{X}$.
Remark 5.28.
If $\min\big{(}\Gamma^{*}(\boldsymbol{\mathbf{v}}),\overline{\Gamma}(\boldsymbol{\mathbf{v}})\big{)}$ is differentiable, taking $\Gamma(\boldsymbol{\mathbf{v}})=(1-\lambda)\min\big{(}\Gamma^{*}(\boldsymbol{\mathbf{v}}),\overline{\Gamma}(\boldsymbol{\mathbf{v}})\big{)}$, with $\lambda\in(0,1)$ small can make $\mathcal{C}$ compact.
Remark 5.29.
When multiple constraints are considered, the stability threshold value $\overline{\Gamma}(\boldsymbol{\mathbf{v}})$ is shared among them. Due to this, only one of the $\Gamma_{i}$ needs to satisfy $\Gamma_{i}(\boldsymbol{\mathbf{v}})\leq\overline{\Gamma}(\boldsymbol{\mathbf{v}})$ as in Theorem 5.26. However, every $\Gamma_{i}$ must respect $\Gamma_{i}(\boldsymbol{\mathbf{v}})\leq\Gamma_{i}^{*}(\boldsymbol{\mathbf{v}})$.
Remark 5.30.
When choosing the initial virtual reference $\boldsymbol{\mathbf{v}}$ in the augmented system (22), it is worth noting that any $\boldsymbol{\mathbf{v}}_{0}\in\mathcal{V}$ such that $(\boldsymbol{\mathbf{x}}_{0},\boldsymbol{\mathbf{v}}_{0})\in\mathcal{C}$ is valid. A desirable choice is
$\boldsymbol{\mathbf{v}}_{0}=\underset{\boldsymbol{\mathbf{v}}\in\mathcal{V}}{\arg\min}~{}V(\boldsymbol{\mathbf{x}}_{0},\boldsymbol{\mathbf{v}}),$ (40)
because it tends to maximize the value of $\Delta$. If the above problem is infeasible, then $\boldsymbol{\mathbf{x}}_{0}$ is not a safe starting condition for the designed DSM.
6 Examples
In this section, we present two nonlinear examples to showcase the presented approach of designing CBFs using dynamic safety margins. For each example, we provide a comparison with candidate CBFs and the ERG. Throughout this section, simulations are computed in MATLAB and optimization problems are solved with MOSEK [34].
6.1 Anthill system
Consider the scalar system
$\dot{x}=(x^{2}-1)x+u,$ (41)
with state space domains $\mathbb{X}=\mathbb{R}$ and $\mathbb{U}=\mathbb{R}$. In the absence of inputs $u=0$, this system asymptotically tends to 00 if $|x_{0}|<1$ and diverges to infinity if $|x_{0}|>1$. Let $\mathcal{U}=[-u_{\max},u_{\max}]$ be the input constraint set, with $u_{\max}\!=\!\frac{4}{3\sqrt{3}}$.
Given the prestabilizing controller
$\pi(x,v)=-(v^{2}-1)v-3xv(x-v),$ (42)
with $v\in\mathbb{V}=\mathbb{R}$, the prestabilized system is
$\dot{x}=f_{\pi}(x,v)=(x-v)^{3}-(x-v),$ (43)
which satisfies the equilibrium conditions $\bar{x}(v)=v$ and $\bar{u}(v)=\pi\big{(}\bar{x}(v),v\big{)}=v-v^{3}$ for all $v\in\mathbb{V}$. The equilibrium point $\bar{x}(v)$ is attractive if $|x_{0}-v|<1$. This can be proven using the Lyapunov function
$V(x,v)=\int_{v}^{x}\!\!\!\!-f_{\pi}(x,v)dx=\frac{1}{2}(x-v)^{2}-\frac{1}{4}(x-v)^{4}.$ (44)
This anthill-shaped function (see Fig. 2) is such that, $\forall v\in\mathbb{V}$,
$V\big{(}\bar{x}(v),v\big{)}$ $=0,$
$V(x,v)$ $>0,\quad\forall x\in\mathcal{D}_{\boldsymbol{\mathbf{v}}}\setminus\{\bar{x}(v)\},$
$\frac{\partial V}{\partial x}f_{\pi}(x,v)$ $=-f_{\pi}^{2}(x,v)<0,\quad\forall x\in\mathcal{D}_{\boldsymbol{\mathbf{v}}}\setminus\{\bar{x}(v)\},$
where $\mathcal{D}_{\boldsymbol{\mathbf{v}}}=(v-1,v+1)$ is the basin of attraction of the equilibrium point $\bar{x}(v)$. The function $V(x,v)$ also satisfies Assumption 4. The stability threshold value (36) is
$\overline{\Gamma}(v)=V(v\pm 1,v)=\frac{1}{4}.$ (45)
To obtain the steady-state admissible set $\mathcal{V}\subset\mathbb{V}$, consider the input constraints at steady-state $|\bar{u}(v)|=|v-v^{3}|\leq u_{\max}$. Using Cardano’s formula to compute the roots of a 3-rd degree polynomial, we obtain $\mathcal{V}=\left[-v_{\max},v_{\max}\right]$, where
$v_{\max}=\frac{(2+\sqrt{3})^{\frac{1}{3}}+(2-\sqrt{3})^{\frac{1}{3}}}{\sqrt{3}}.$ (46)
The input constraints also induce state constraints for the prestabilized system. By setting $\pi(x,v)=\pm u_{\max}$ and solving for $x$, we obtain the reference-dependent state constraint set.
$\mathcal{X}_{v}=\begin{cases}\emptyset,\quad&v<-4/\sqrt{3},\\ \frac{1}{2}v+\left[\delta^{-}(v),-\delta^{-}(v)\right],\quad&v\in\left[-4/\sqrt{3},0\right),\\ \mathbb{R},\quad&v=0,\\ \frac{1}{2}v+\left[-\delta^{+}(v),\delta^{+}(v)\right],\quad&v\in\left(0,4/\sqrt{3}\right],\\ \emptyset,\quad&v>4/\sqrt{3},\end{cases}$
where
$\delta^{\pm}(v)=\frac{\sqrt{-3v^{4}+12v(v\pm u_{\max})}}{6v}.$ (47)
The resulting threshold value is
$\Gamma^{*}(v)\!=\!\begin{cases}-\infty,\;&v\in(-\infty,-1-v_{\max}],\\ -V\left(\frac{1}{2}v+\delta^{-}(v),v\right),\;&v\in(-1-v_{\max},-v_{\max}],\\ V\left(\frac{1}{2}v+\delta^{-}(v),v\right),\;&v\in(-v_{\max},1-v_{\max}),\\ \infty,\;&v\in[1-v_{\max},v_{\max}-1],\\ V\left(\frac{1}{2}v+\delta^{+}(v),v\right),\;&v\in(v_{\max}-1,v_{\max}),\\ -V\left(\frac{1}{2}v+\delta^{+}(v),v\right),\;&v\in[v_{\max},v_{\max}+1),\\ -\infty,\;&v\in[v_{\max}+1,\infty).\end{cases}$
Noting that $V\big{(}\frac{1}{2}v+\delta^{+}(v),v\big{)}=\frac{1}{4}=\overline{\Gamma}(v)$ if $v=v_{\max}\pm 1$, it can be shown that
$\Gamma(v)=0.99\max\left(\min\left(\Gamma^{*}(v),\overline{\Gamma}(v)\right),-\overline{\Gamma}(v)\right),$ (48)
is a continuously differentiable function satisfying the conditions of Theorem 5.26. Thus, $\Delta(x,v)=\Gamma(v)-V(x,v)$ is a Lyapunov-based DSM for $\pi$ and a CBF by Theorem 3.7. Furthermore, the set $\mathcal{C}\subset\mathbb{X}\times\mathcal{V}$ is compact in $\mathbb{X}\times\mathbb{V}$ (see Remark 5.28).
Figure 2: Anthill-shaped Lyapunov function $V(x,v)$ at $v=0.5$ and the associated threshold value $\Gamma^{*}(v)$ and stability threshold value $\overline{\Gamma}(v)$.
We now consider the state constraints $\mathcal{X}=[-x_{\max},x_{\max}]$, with $x_{\max}=2$. Since $x_{\max}>v_{\max}$ and (41) is a first order system (no overshoot), these constraints have already been enforced by the restriction $v\in\mathcal{V}$ induced by $\mathcal{C}$. Thus, no further action is needed.
Remark 6.31.
It is interesting to note that, given $x_{\max}>v_{\max}$, the value of $x_{\max}$ plays no role in the DSM-based CBF. Indeed, the primary concern is to prevent instability by maintaining control authority over the system: any loss in control authority makes constraint violation inevitable, regardless of the specific value of $x_{\max}$.
Given the initial conditions $x_{0}=0.51$, $v_{0}=0.51$, the desired reference $r=1.5$, and the nominal control law $\kappa(x,r)=\pi(x,r)$, we consider the following constrained control strategies:
  1. (a)
    DSM-based CBF: We use $\eta=0.01$, the class $\mathscr{K}_{\infty}$ function $\alpha:c\mapsto 1.8c$, and $\rho:(v,r)\mapsto r-v$.
  2. (b)
    ERG: We use $\rho_{\text{ERG}}:(v,r)\mapsto 100(r-v)$.
  3. (c)
    Candidate CBF: We use the functions $h_{1}(x)=x_{\max}-x$ and $h_{2}(x)=x_{\max}+x$ and tune their class $\mathscr{K}_{\infty}$ functions to illustrate the tradeoff between performance and feasibility that candidate CBFs suffer from.
The simulation results for each approach are shown in Fig. 3. Both the ERG and the proposed DSM-CBF successfully enforce safety and reach the closest safe point to the desired reference but, the DSM-CBF exhibits performance that more closely resembles the nominal behavior. With enough tuning, the candidate CBFs were able to also maintain safety for the fist 10 seconds, albeit with poor performance and no guarantee they will remain feasible afterwards. This is unsurprising since the candidate CBFs were constructed without any understanding of the inherent control authority constraint. Conversely, the DSM-based CBF provided a systematic approach that successfully identified the constraints induced by the stability region of the nominal or prestabilizing controllers.
Figure 3: Closed-loop behavior for the anthill system using each of the considered constrained control approaches. The dashed green and blue lines represent the evolution of the virtual reference $v(t)$ for the ERG and the DSM-CBF approaches, respectively. Note that $v(t)$ may decrease in our approach, which is not possible in the ERG. This feature enlarges the feasible region of the CBF-based QP and allows the safe input $\beta(x,v)$ to be closer to the nominal input $\pi(x,r)$. In addition, we show the performance of the candidate CBFs for different class $\mathscr{K}_{\infty}$ functions $\alpha_{1}=\alpha_{2}=\alpha:c\mapsto a_{i}c$, where $a_{i}\in\{7,0.25,0.15,0.1,0.07\}$. Larger $a_{i}$ corresponds to better input tracking, but also an earlier failure due to infeasibility, marked by a magenta ‘$\times$’ marker.
6.2 Overhead Crane
Consider the dynamics of an overhead crane [35]
$M(\boldsymbol{\mathbf{q}})\ddot{\boldsymbol{\mathbf{q}}}+V_{m}(\boldsymbol{\mathbf{q}},\dot{\boldsymbol{\mathbf{q}}})\dot{\boldsymbol{\mathbf{q}}}+G(\boldsymbol{\mathbf{q}})=Bu,$ (49)
where the degrees of freedom $\boldsymbol{\mathbf{q}}=[x;~{}\theta]$ are the gantry position $x$ and the payload angle $\theta$, and
$M(\boldsymbol{\mathbf{q}})$ $=\begin{bmatrix}m_{c}+m_{p}&-m_{p}L\cos\theta\\ -m_{p}L\cos\theta&m_{p}L^{2}\end{bmatrix},\quad B=\begin{bmatrix}1\\ 0\end{bmatrix},$
$V_{m}(\boldsymbol{\mathbf{q}},\dot{\boldsymbol{\mathbf{q}}})$ $=\begin{bmatrix}0&m_{p}L\dot{\theta}\sin\theta\\ 0&0\end{bmatrix},\quad G(\boldsymbol{\mathbf{q}})=\begin{bmatrix}0\\ m_{p}gL\sin\theta\end{bmatrix},$
where $m_{c},m_{p}>0$ represent the gantry and payload masses, respectively, $L>0$ is the length of the rod connecting the crane and payload, and $g=9.81$ is the acceleration of gravity (see Fig. 4).
Figure 4: Overhead crane system from [35].
It can be shown that, letting $\boldsymbol{\mathbf{x}}=[\boldsymbol{\mathbf{q}};~{}\dot{\boldsymbol{\mathbf{q}}}]$, the system is control-affine and can be written as in (1). We consider the prestabilizing PD control law given in [35]
$\pi(\boldsymbol{\mathbf{x}},v)=-k_{p}(x-v)-k_{d}\dot{x},$ (50)
where $k_{p}>0$, $k_{d}>0$ are the proportional and derivative gains. The equilibrium mapping is $\bar{x}:v\mapsto[v~{};0~{};0~{};0]$ and the reference-dependent Lyapunov function is
$V(\boldsymbol{\mathbf{x}},v)=\frac{1}{2}\dot{\boldsymbol{\mathbf{q}}}^{\top}M(\boldsymbol{\mathbf{q}})\dot{\boldsymbol{\mathbf{q}}}+m_{p}gL(1-\cos\theta)+\frac{1}{2}k_{p}(x-v)^{2},$
[…]
  1. 1.
    Position constraints: $x_{\text{min}}\leq x\leq x_{\text{max}}$. For each of the two bounds, we take
    $\Gamma_{1}(v)$ $=\Gamma_{1}^{*}(v)=\frac{1}{2}k_{p}(v-x_{\text{min}})|v-x_{\text{min}}|,$
    $\Gamma_{2}(v)$ $=\Gamma_{2}^{*}(v)=\frac{1}{2}k_{p}(x_{\text{max}}-v)|x_{\text{max}}-v|.$
  2. 2.
    Input constraints: $|u|\leq u_{\text{max}}$, with $u_{\text{max}}>0$. Solving (32) yields the constant function
    $\Gamma_{3}(v)=\Gamma_{3}^{*}(v)=\frac{m_{c}u_{\text{max}}^{2}}{2(m_{c}k_{p}+k_{d}^{2})}.$
  3. 3.
    Angle constraints: $|\theta|\leq\theta_{\text{max}}$, with $\theta_{\text{max}}\in(0,\pi/2)$. Solving (32) yields the constant function
    $\Gamma_{4}(v)=\Gamma_{4}^{*}(v)=m_{p}gL(1-\cos\theta_{\text{max}}).$
    We note that this threshold value is strictly smaller than the stability threshold value (36), i.e., $\overline{\Gamma}(v)=2m_{p}gL$. As such, enforcing the angle constraints ensures $\boldsymbol{\mathbf{x}}\in\mathcal{D}_{v}$.
  4. 4.
    Payload constraints: $x+L\sin\theta\leq p_{\text{max}}$, with $p_{\text{max}}\geq x_{\text{max}}$. For this constraint, we were unable to find an analytical expression for $\Gamma_{5}^{*}(v)$ associated to $V$. However, as detailed in [28, Proposition 4], we can define
    $\underline{V}(\boldsymbol{\mathbf{x}},v)=\frac{1}{2}\dot{\boldsymbol{\mathbf{q}}}^{\top}M(\boldsymbol{\mathbf{x}})\dot{\boldsymbol{\mathbf{q}}}+\frac{4}{\pi^{2}}m_{p}gL\theta^{2}+\frac{1}{2}k_{p}(x-v)^{2},$
    which satisfies $\underline{V}(\boldsymbol{\mathbf{x}},v)\leq V(\boldsymbol{\mathbf{x}},v)$, $\forall\theta\in(-\pi/2,\pi/2)$. Moreover, since $x_{\text{max}}\leq p_{\text{max}}$, we can perform a planar embedding of the payload constraint
    $x+L\theta\leq p_{\text{max}}\quad\implies\quad x+L\sin\theta\leq p_{\text{max}},$ (51)
    which holds $\forall x\leq x_{\text{max}}$ and $\forall\theta\in[-\pi/2,\pi/2]$. With these approximations, the following function satisfies the condition of Theorem 5.26:
    $\Gamma_{5}(v)=\frac{4k_{p}m_{p}g}{8m_{p}g+Lk_{p}\pi^{2}}(p_{\text{max}}-v)|p_{\text{max}}-v|.$
The resulting DSMs are $\Delta_{i}(\boldsymbol{\mathbf{x}},v)=\Gamma_{i}(v)-V(\boldsymbol{\mathbf{x}},v)$ for all $i\in\{1,\ldots,5\}$. By Theorem 4.22, these DSMs have the control-sharing property. For all simulations, we use linear class $\mathscr{K}_{\infty}$ functions given by $\alpha_{i}:c\mapsto\alpha_{i}c$ where $\alpha_{i}>0$ represents both the function and the scalar gain despite the slight abuse of notation. We choose $\alpha_{1}=\alpha_{2}=10$, $\alpha_{3}=20$, $\alpha_{4}=80$ and $\alpha_{5}=2.2$. The navigation field we use is $\rho(v,r)=r-v$.
The following approaches are used for comparison.
  1. (a)
    Nominal: The nominal controller is a PD control law $\kappa(\boldsymbol{\mathbf{x}},r)=-k_{p,\kappa}(x-r)-k_{d,\kappa}\dot{x}$, where the gains $k_{p,\kappa}>0$ and $k_{d,\kappa}>0$ can differ from the ones in (50) and $r\in\mathbb{R}$ is the target reference.
  2. (b)
    ERG: The explicit reference governor uses the same dynamic safety margins as our approach, but changing $k_{p}\to k_{p,\kappa}$ and $k_{d}\to k_{d,\kappa}$ in all the expressions. The navigation field is $\rho_{\text{ERG}}(v,r)=1000(r-v)$. The ERG-based, closed-loop system is
    $\begin{bmatrix}\dot{\boldsymbol{\mathbf{x}}}\\ \dot{v}\end{bmatrix}=\begin{bmatrix}f(\boldsymbol{\mathbf{x}})+g(\boldsymbol{\mathbf{x}})\kappa(\boldsymbol{\mathbf{x}},v)\\ \min_{i\in\{1,\ldots,5\}}\Delta_{i}(\boldsymbol{\mathbf{x}},v)\rho_{\text{ERG}}(v,r)\end{bmatrix}.$ (52)
  3. (c)
    Candidate CBF: For each constraint, we design a candidate CBF.
    1. 1)
      Position constraints: $h_{1}(\boldsymbol{\mathbf{x}})=\gamma_{1}(x-x_{\text{min}})+\dot{x}$ for $x_{\text{min}}\leq x$ and $h_{2}(\boldsymbol{\mathbf{x}})=\gamma_{2}(x_{\text{max}}-x)-\dot{x}$ for $x\leq x_{\text{max}}$.
    2. 2)
      Input constraints are added as upper and lower bounds for the optimization variable $-u_{\text{max}}\leq u\leq u_{\text{max}}$.
    3. 3)
      Angle constraints: $h_{3}(\boldsymbol{\mathbf{x}})=\gamma_{3}(\theta_{\text{max}}+\theta)+\dot{\theta}$ for $-\theta_{\text{max}}\leq\theta$ and $h_{4}(\boldsymbol{\mathbf{x}})=\gamma_{4}(\theta_{\text{max}}-\theta)-\dot{\theta}$ for $\theta\leq\theta_{\text{max}}$.
    4. 4)
      Payload constraints: $h_{5}(\boldsymbol{\mathbf{x}})=\gamma_{5}(p_{\text{max}}-x-L\sin\theta)-\dot{x}-L\dot{\theta}\cos\theta$ for $x+L\sin\theta\leq p_{\text{max}}$.
    Where $\gamma_{i}$ are positive constants. For all simulations, we use $\gamma_{i}=6$ and $\tilde{\alpha}_{i}=8,\forall i\in\{1,\ldots,4\}$, and $\gamma_{5}=4$, $\tilde{\alpha}_{5}=3.5$, where $\tilde{\alpha}_{i}:c\mapsto\tilde{\alpha}_{i}c$ is the class $\mathcal{K}_{\infty}$ function associated with $h_{i}$. It must be noted that all $h_{i}$ described above are candidate CBFs, meaning that they may not satisfy Definition 2 or have the control-sharing property given in Definition 4.21.
Table 1: Overhead Crane Simulation Parameters
$m_{c}$ (kg) $1$ $k_{p}$ $1$ $x_{\text{min}}$ (m) $1.1$ $p_{\text{max}}$ (m) $1.1$
$m_{p}$ (kg) $0.5$ $k_{d}$ 0.1 $x_{\text{max}}$ (m) $1.1$ $\boldsymbol{\mathbf{x}}(0)$ (-) 00
$L$ (m) $0.7$ $k_{p,\kappa}$ $10$ $u_{\text{max}}$ (N) $4$ $v(0)$ (m) $0.1$
$\eta$ (-) 0.01 $k_{d,\kappa}$ $4$ $\theta_{\text{max}}$ (deg) $10$ $r$ (m) $1$
General parameters for the example are provided in Table 1. For the first simulation, shown in Fig. 5, we relax the angle constraint to $|\theta|\leq\pi/2$, which effectively makes it inactive. In this case, all constrained control approaches achieve the safety objective. In terms of performance, the Candidate CBF approach tracks the nominal behavior better than the proposed DSM-CBF approach, whereas the ERG exhibits the slowest response. The second simulation, shown in Fig. 6, includes the angle constraint $|\theta|\leq\theta_{\text{max}}$. The ERG and our proposed DSM-CBF approaches successfully enforce the constraints, with the proposed method achieving better performance. However, the Candidate CBF-based controller becomes infeasible despite our best efforts to tune $\gamma_{i}$ and $\tilde{\alpha}_{i}$. This is unsurprising since candidate CBFs do not guarantee persistent feasibility.
Figure 5: Closed-loop behavior in the absence of angle constraints. Even though the nominal controller violates the constraints, the Candidate CBF, ERG, and proposed DSM-CBF approaches all successfully enforce the safety requirements.
Figure 6: Closed-loop behavior in the presence of angle constraints. The ‘$\times$’ marks the instant in which the Candidate CBF policy becomes infeasible, likely due to the unavoidable backswing experienced by the payload angle $\theta(t)$. Conversely, the ERG and proposed DSM-CBF approaches successfully enforce the safety requirements.
7 Conclusion
This paper showed how dynamic safety margins can be used to systematically construct control barrier functions with the control-sharing property. The work also highlighted a Lyapunov-based approach to compute DSM-CBFs and presented simulation results that demonstrate their ability to guarantee that the CBF optimization problem is persistently feasible for systems with high relative degree, multiple state and input constraints, and that are vulnerable to a loss of control authority. Future work includes continuing to explore connections between the CBF and reference governor literatures and robustifying the presented results to account for bounded disturbances.