### v3

# Using Dynamic Safety Margins as Control Barrier Functions

Abstract
This paper presents an approach to design control barrier functions (CBFs) for arbitrary state and input constraints using tools from the reference governor literature. In particular, it is shown that dynamic safety margins (DSMs) are CBFs for an augmented system obtained by concatenating the state with a virtual reference. The proposed approach is agnostic to the relative degree and can handle multiple state and input constraints using the control-sharing property of CBFs. The construction of CBFs using Lyapunov-based DSMs is then investigated in further detail. Numerical simulations show that the method outperforms existing DSM-based approaches, while also guaranteeing safety and persistent feasibility of the associated optimization program.
{IEEEkeywords}
Constrained control, Control barrier functions, Predictive control for nonlinear systems.
1 Introduction
[…]
Control barrier functions (CBFs) have become a popular tool for deriving optimization-based constrained control laws that are safe, easy to implement, and achieve good performance [1, 2, 3]. Despite their success, the widespread use of CBFs is limited by the absence of a systematic method to synthesize them for general classes of systems with arbitrary state and input constraints.
Given a state constraint set with a smooth boundary, it is usually straightforward to find a differentiable function of the state vector that is positive for constraint-admissible states, and negative otherwise. We call such functions candidate CBFs. They are commonly used in practice to design CBF-based controllers and, with tuning, can achieve great performance (e.g., [2]). However, candidate CBFs lack many safety guarantees that valid CBFs provide. In particular, the CBF-based optimization program governing the control law may become infeasible when using candidate CBFs, which is unacceptable for safety-critical systems. Ultimately, CBFs are certificates of control invariance. Therefore, if the original state constraint set is not control invariant, it is impossible to find a valid CBF associated to it. In these cases, we must resort to finding CBFs for a, hopefully large, control invariant subset of the state constraint set. In this work, we use tools from the reference governor literature [4] to systematically design valid CBFs given arbitrary state and input constraints.
Figure 1: Comparison of CBF-based control (top), which filters a nominal input signal $\boldsymbol{\mathbf{u}}_{n}$ to obtain a safe input $\boldsymbol{\mathbf{u}}^{*}$, with a specific reference governor approach, the command governor (CG). The CG filters the desired reference signal $\boldsymbol{\mathbf{r}}$ to obtain a safe virtual reference $\boldsymbol{\mathbf{v}}^{*}$ for the prestabilizing controller.
1.1 Related Works
While the term control barrier function can be traced back to the early 2000’s (see [5] for a historical account), the modern definition that we consider in this work was first presented in [6] under the name of zeroing CBFs. Since then, many works have studied different properties of CBFs and the systems they are used in: [7, 8] address the relative degree of CBFs; [9] defines control-sharing CBFs to enforce multiple constraints at once; [10, 11] consider discrete-time and sampled-data systems, respectively; [12, 13] study robust CBFs.
Most relevant to the present manuscript, we review some works that study the design of CBFs and feasibility of the CBF-based optimization program. In [14], sum of squares (SOS) techniques are used to synthesize CBFs for semi-algebraic sets and polynomial systems using an alternating descent algorithm. In [15], the authors use SOS techniques to synthesize polynomial control Lyapunov functions, and then present an extension to construct CBFs. In [16], the authors introduce a variable modifying the decay rate of the CBF to guarantee feasibility of the CBF-based program when the state is in the interior of the safe set even under input constraints. However, their approach provides no feasibility guarantee when the state is on the boundary of the safe set. In [17], the authors use control Lyapunov functions (CLFs) to design a continuum of valid CBFs and use tools from hybrid systems theory to study the resulting safety properties. In [18], the authors design CBFs for Euler–Lagrange systems. While the results are promising and the class of systems is relevant, the approach is limited to box constraints. In [19], the authors learn CBFs from human demonstrations but they provide no guarantees that the learned CBFs are valid. In [20], given a valid CBF, the authors use learning to construct a CBF for a larger control invariant set using a combination of safe and unsafe trajectories of the system. Nonetheless, finding the initial, valid CBF remains an open question. In [21], the authors use a backup control policy to enlarge a small (but easy to find) control invariant set. The enlarged control invariant set is then used to derive a backup CBF. These backup CBFs rarely have a closed form, which makes them difficult to implement. In [22], the authors use maximal output admissible sets (MOASs) to design discrete-time CBFs for arbitrary state and input constraints. While the approach works well for linear systems, finding the MOAS for nonlinear systems remains an open question.
This work expands the underlying theme of [21, 22], where CBFs are obtained starting from a prestabilizing (or backup) controller. To do this, we adopt the notion of dynamic safety margin (DSM) from the explicit reference governor (ERG) framework [23]. Given a fixed reference, a DSM measures the distance to constraint violation of a prestabilized system trajectory as it converges to the associated equilibrium point. In [24], the authors define governor-parameterized barrier functions (PBFs), which are based on CBFs, and show that they are DSMs. While the results are limited to linear systems, their work is one of the few connections between the reference governor and CBF literatures. In contrast to their result, we show that DSMs are CBFs for an augmented system consisting of the concatenated state and reference vectors. These works serve as motivation for exploring the connections between reference governor approaches like the command governor (CG) [25], which filters a nominal reference, and CBF-based control, which filters a nominal input. Both approaches ultimately modify a nominal, but potentially unsafe signal to guarantee closed-loop safety (see Fig. 1).
1.2 Contributions
The contributions of this manuscript are:
  • We leverage the control-sharing property presented in [9] to define a natural generalization of control barrier functions (CBFs) to (vector-valued) CBFs.
  • We show that dynamic safety margins (DSMs) are (vector-valued) CBFs for an augmented system consisting of the concatenation of the state vector and the reference of the prestabilizing controller used to construct the DSM.
  • We show that the DSM-CBFs support constraints with arbitrary relative degree.
  • We formulate the DSM-CBF optimization-based policy and show that it is safe and persistently feasible. Moreover, the optimization problem is a quadratic program (QP) when the input constraints are polyhedral.
  • We provide conditions that guarantee the DSM-CBF optimization-based policy is locally Lipschitz continuous. This guarantees existence and uniqueness of the closed-loop trajectory.
  • We extend the Lyapunov-based DSMs presented in [23] to inadmissible references. This extension is needed to use them as DSM-CBFs. Moreover, we show that any class $\mathscr{K}_{\infty}$ function can be used with Lyapunov-based DSM-CBFs.
  • We present a comparison of our proposed approach with the explicit reference governor (ERG) [23], candidate CBFs and backup CBFs [21] in two nonlinear systems. The numerical simulations show that the DSM-CBFs we propose outperform the ERG, while boasting strict safety guarantees that are otherwise lost when candidate CBF-based control becomes infeasible. The examples also show that, while backup CBFs can achieve better performance than DSM-CBFs, their implementation is challenging in the presence of multiple constraints or loss of control authority.
The rest of the paper is organized as follows: Section 2 provides formal definitions for control invariance, CBFs, and DSMs; Section 3 proves that DSMs are CBFs, shows the relative degree of the constructed CBF is irrelevant for safety, formulates the CBF-based optimization problem, which is a quadratic program (QP) for polyhedral input constraints, and details some conditions that guarantee the DSM-CBF policy is locally Lipschitz continuous; Section 4 provides a systematic approach based on Lyapunov functions to construct DSMs and highlights some desirable properties that such DSMs have; Section 5 summarizes the presented CBF synthesis approach and contextualizes it with respect to the existing reference governor literature; Section 6 presents the numerical simulations and comparisons; Section 7 concludes the paper.
1.3 Notation
Throughout the paper, only vector quantities are presented in bold font: $\boldsymbol{\mathbf{x}}\in\mathbb{R}^{n}$. The notation $[\boldsymbol{\mathbf{x}};~{}\boldsymbol{\mathbf{y}}]$ denotes the vertical stacking of two vectors $\boldsymbol{\mathbf{x}}$ and $\boldsymbol{\mathbf{y}}$ into a column vector. A continuous function $\alpha:[0,\infty)\to[0,\infty)$ is said to belong to class $\mathscr{K}_{\infty}$ if it is strictly increasing, $\alpha(0)=0$, and $\lim_{r\to\infty}\alpha(r)=\infty$. The symbol $\subset$ denotes set inclusion, but not necessarily strict inclusion. Given a set $\mathcal{S}$, we denote by $\operatorname{Int}\left(\mathcal{S}\right)$ the interior of $\mathcal{S}$ and by $\partial\mathcal{S}$ the boundary of $\mathcal{S}$ (i.e. $\partial\mathcal{S}=\text{closure}(\mathcal{S})\setminus\operatorname{Int}\left(\mathcal{S}\right)$). Given a subset $\mathcal{C}\subset\mathbb{R}^{n}$ of the state space $\mathbb{R}^{n}$, we generally denote by $\tilde{\mathcal{C}}\subset\mathbb{R}^{n}\times\mathbb{R}^{l}$ its counterpart in the augmented state-reference space. Let $h:\mathcal{D}\to\mathbb{R}$ be continuously differentiable. We say $\boldsymbol{\mathbf{x}}\in\mathcal{D}$ is a regular point of $h$ if $\partial h(\boldsymbol{\mathbf{x}})/\partial\boldsymbol{\mathbf{x}}\neq 0$. We say $c\in\mathbb{R}$ is a regular value of $h$ if every $\boldsymbol{\mathbf{x}}\in\mathcal{D}$ such that $h(\boldsymbol{\mathbf{x}})=c$ is a regular point.
2 Preliminaries
This section recalls the notion of control invariance, control barrier functions, and dynamic safety margins. Consider a control-affine system
$\dot{\boldsymbol{\mathbf{x}}}=f(\boldsymbol{\mathbf{x}})+g(\boldsymbol{\mathbf{x}})\boldsymbol{\mathbf{u}},$ (1)
where $f:\mathbb{R}^{n}\to\mathbb{R}^{n}$ and $g:\mathbb{R}^{n}\to\mathbb{R}^{n\times m}$ are locally Lipschitz continuous functions. The system is subject to state $\boldsymbol{\mathbf{x}}\in\mathcal{X}$ and input $\boldsymbol{\mathbf{u}}\in\mathcal{U}$ constraints, where $\mathcal{X}\subset\mathbb{R}^{n}$ and $\mathcal{U}\subset\mathbb{R}^{m}$ are closed sets.
2.1 Control Invariance
Control invariance is a desirable property of state-space subsets that is closely related to safety. Roughly, a set is control invariant if there exists a suitable choice of inputs such that the trajectory of the closed-loop system remains in the set.
Definition 1 ([26])
A set $\mathcal{C}\subset\mathbb{R}^{n}$ is control invariant if, $\forall\boldsymbol{\mathbf{x}}_{0}\in\mathcal{C}$, there exists an admissible input signal $\boldsymbol{\mathbf{u}}:[0,\infty)\to\mathcal{U}$ such that at least one of the trajectories of the closed-loop system (1), with $\boldsymbol{\mathbf{x}}(0)=\boldsymbol{\mathbf{x}}_{0}$, satisfies $\boldsymbol{\mathbf{x}}(t)\in\mathcal{C}$ for all $t$ in its interval of existence $\mathcal{J}(\boldsymbol{\mathbf{x}}_{0})\subset[0,\infty)$.
Note that this concept is sometimes called weak control invariance [26] because it doesn’t require existence and uniqueness of the trajectory. Nonetheless, it will simplify the presentation of the upcoming results. Assumptions for the existence and uniqueness of solutions will be presented separately. Central to the upcoming discussion is Nagumo’s theorem characterizing control invariance in terms of Bouligand’s tangent cone.
Theorem 1 (Nagumo [27])
A compact set $\mathcal{C}\subset\mathbb{R}^{n}$ is control invariant if and only if
$\forall\boldsymbol{\mathbf{x}}\in\mathcal{C},\quad\exists\boldsymbol{\mathbf{u}}\in\mathcal{U},\quad f(\boldsymbol{\mathbf{x}})+g(\boldsymbol{\mathbf{x}})\boldsymbol{\mathbf{u}}\in\mathcal{T}_{\mathcal{C}}(\boldsymbol{\mathbf{x}}),$ (2)
where $\mathcal{T}_{\mathcal{C}}(\boldsymbol{\mathbf{x}})$ is the tangent cone to $\mathcal{C}$ in $\boldsymbol{\mathbf{x}}$ [26, Definition 3.1].
While this characterization was groundbreaking, working with the tangent cone $\mathcal{T}_{\mathcal{C}}(\boldsymbol{\mathbf{x}})$ is cumbersome. Fortunately, when the compact set $\mathcal{C}$ has a smooth boundary (i.e., can be represented as the superlevel set of a continuously differentiable function), the tangent cone on the boundary is reduced to the tangent half-space shifted to the origin and we have the following characterization.
Theorem 2 ([28, Proposition 4.3.7])
Let $h:\mathcal{D}\to\mathbb{R}$ be a continuously differentiable function with regular value 00 and suppose $\mathcal{C}=\{\boldsymbol{\mathbf{x}}\in\mathcal{D}:h(\boldsymbol{\mathbf{x}})\geq 0\}$ is compact. The set $\mathcal{C}$ is control invariant if and only if
$\forall\boldsymbol{\mathbf{x}}\in\partial\mathcal{C},\quad\exists\boldsymbol{\mathbf{u}}\in\mathcal{U},\quad L_{f}h(\boldsymbol{\mathbf{x}})+L_{g}h(\boldsymbol{\mathbf{x}})\boldsymbol{\mathbf{u}}\geq 0,$ (3)
where $L_{f}h$ and $L_{g}h$ are the Lie derivatives of $h$ along $f$ and $g$, respectively.
Note that [28, Proposition 4.3.7] is more general than Theorem 2: it allows the set $\mathcal{C}$ to be defined using multiple continuously differentiable functions. We use this fact in upcoming results.
Although much simpler than evaluating the tangent cone, the conditions of Theorem 2 apply only on the boundary of the set $\partial\mathcal{C}$, which makes implementing them difficult in practice. This motivated the development of control barrier functions, which allow us to extend these conditions to the entirety of the set $\mathcal{C}$.
2.2 Control Barrier Functions
Control barrier functions are real-valued functions that certify the control invariance of their superlevel set.
Definition 2 ([6])
A continuously differentiable function $h:\mathcal{D}\to\mathbb{R}$ is a control barrier function (CBF) if there exists a class $\mathscr{K}_{\infty}$ function $\alpha:[0,\infty)\to[0,\infty)$ such that
$\forall\boldsymbol{\mathbf{x}}\in\mathcal{C},\quad\sup_{\boldsymbol{\mathbf{u}}\in\mathcal{U}}\left[L_{f}h(\boldsymbol{\mathbf{x}})+L_{g}h(\boldsymbol{\mathbf{x}})\boldsymbol{\mathbf{u}}\right]\geq-\alpha\big{(}h(\boldsymbol{\mathbf{x}})\big{)},$ (4)
where
$\mathcal{C}=\{\boldsymbol{\mathbf{x}}\in\mathcal{D}\mid h(\boldsymbol{\mathbf{x}})\geq 0\}.$ (5)
With this definition, the conditions of Theorem 2 were extended to the entire set $\mathcal{C}$, as shown in the next result.
Lemma 1
Assume $\mathcal{U}\!\subset\!\mathbb{R}^{m}$ is compact. Let $\mathcal{C}\!\subset\!\mathbb{R}^{n}$, defined as in (5), be compact and let 00 be a regular value of $h$. The set $\mathcal{C}$ is control invariant if and only if $h$ is a CBF.
Proof 2.3.
Although the proof is based on [6], they showed this result for barrier functions, not control barrier functions. Let $\mathcal{C}\subset\mathbb{R}^{n}$ be control invariant and define the set-valued map $C(c)=\{\boldsymbol{\mathbf{x}}\in\mathcal{D}\mid 0\leq h(\boldsymbol{\mathbf{x}})\leq c\}$. Note that for any $c\geq 0$, $C(c)\subset\mathcal{C}$ is compact in $\mathbb{R}^{n}$. Next, define the function $\hat{\alpha}:[0,\infty)\to\mathbb{R}$
$\hat{\alpha}(c)=-\inf_{\boldsymbol{\mathbf{x}}\in C(c)}\sup_{\boldsymbol{\mathbf{u}}\in\mathcal{U}}\left[L_{f}h(\boldsymbol{\mathbf{x}})+L_{g}h(\boldsymbol{\mathbf{x}})\boldsymbol{\mathbf{u}}\right].$ (6)
Since $\mathcal{U}$ is compact, it follows by the extreme value theorem that $\xi:\mathcal{D}\to\mathbb{R}$, defined as
$\xi(\boldsymbol{\mathbf{x}})=\sup_{\boldsymbol{\mathbf{u}}\in\mathcal{U}}\left[L_{f}h(\boldsymbol{\mathbf{x}})+L_{g}h(\boldsymbol{\mathbf{x}})\boldsymbol{\mathbf{u}}\right],$ (7)
is continuous on $\mathcal{D}$. Moreover, it can be shown that the set-valued map $C$ is both uppper- and lower-hemicontinuous. Thus, it follows by Berge’s maximum theorem [29], that the value function $\hat{\alpha}$ is continuous on $[0,\infty)$. By construction, $\hat{\alpha}$ is non-decreasing. By Theorem 2, for any $\boldsymbol{\mathbf{x}}\in\partial\mathcal{C}$, $\exists\boldsymbol{\mathbf{u}}\in\mathcal{U}$ such that $L_{f}h(\boldsymbol{\mathbf{x}})+L_{g}h(\boldsymbol{\mathbf{x}})\boldsymbol{\mathbf{u}}\geq 0$. Thus, $\forall\boldsymbol{\mathbf{x}}\in\partial\mathcal{C}=\mathcal{C}(0)$, $\xi(\boldsymbol{\mathbf{x}})\geq 0$. As a result,
$\hat{\alpha}(0)=-\inf_{\boldsymbol{\mathbf{x}}\in\partial\mathcal{C}}\xi(\boldsymbol{\mathbf{x}})\leq 0.$ (8)
There always exists a class $\mathscr{K}_{\infty}$ function $\alpha:[0,\infty)\to[0,\infty)$ that upper-bounds $\hat{\alpha}$. Finally, let $\boldsymbol{\mathbf{x}}\in\mathcal{C}$ be given and note
$\xi(\boldsymbol{\mathbf{x}})\geq\inf_{\boldsymbol{\mathbf{x}}_{0}\in\mathcal{C}(h(\boldsymbol{\mathbf{x}}))}\xi(\boldsymbol{\mathbf{x}}_{0})=-\hat{\alpha}\big{(}h(\boldsymbol{\mathbf{x}})\big{)}\geq-\alpha\big{(}h(\boldsymbol{\mathbf{x}})\big{)}.$
So, we conclude that $h$ is a CBF.
On the other hand, assume $h$ is a CBF, let $\alpha\in\mathscr{K}_{\infty}$ be such that (4) holds, and let $\boldsymbol{\mathbf{x}}\in\partial\mathcal{C}$ be given. By compactness of $\mathcal{U}$, there exists $\boldsymbol{\mathbf{u}}\in\mathcal{U}$ such that $L_{f}h(\boldsymbol{\mathbf{x}})+L_{g}h(\boldsymbol{\mathbf{x}})\boldsymbol{\mathbf{u}}\geq-\alpha\big{(}h(\boldsymbol{\mathbf{x}})\big{)}=0$, where the last equality holds because $h(\boldsymbol{\mathbf{x}})=0$. Since $\mathcal{C}$ is compact and 00 is a regular value of $h$, we can apply Theorem 2 to conclude $\mathcal{C}$ is control invariant.
Notice compactness of $\mathcal{U}$ implies $\xi:\mathcal{D}\to\mathbb{R}$ is continuous, and therefore $\hat{\alpha}:[0,\infty)\to\mathbb{R}$. Without it, we lose necessity in Lemma 1. However, the sufficiency in Lemma 1 (i.e., the CBF certifies control invariance of its compact superlevel set) is maintained even when the compactness of $\mathcal{U}$ is exchanged for the less restrictive condition
$\forall\boldsymbol{\mathbf{x}}\in\mathcal{C},\quad\mathcal{K}_{\alpha}(\boldsymbol{\mathbf{x}})\neq\emptyset,$ (9)
where
$\mathcal{K}_{\alpha}(\boldsymbol{\mathbf{x}})\!=\!\{\boldsymbol{\mathbf{u}}\in\mathcal{U}\mid L_{f}h(\boldsymbol{\mathbf{x}})+L_{g}h(\boldsymbol{\mathbf{x}})\boldsymbol{\mathbf{u}}+\alpha\big{(}h(\boldsymbol{\mathbf{x}})\big{)}\!\geq\!0\}.$ (10)
In this work we drop the compactness requirement on $\mathcal{U}$ because only the sufficiency statement in Lemma 1 is relevant and we will design CBFs for which (9) holds.
CBFs can be used to design add-on modules that enforce constraint satisfaction by filtering the control input. Specifically, let $h$ be a CBF such that $\mathcal{C}\subset\mathcal{X}$ and consider the closed-loop system
$\dot{\boldsymbol{\mathbf{x}}}=f(\boldsymbol{\mathbf{x}})+g(\boldsymbol{\mathbf{x}})\beta(\boldsymbol{\mathbf{x}}),$ (11)
where $\beta:\mathcal{C}\to\mathcal{U}$ is the optimization-based control policy
$\beta(\boldsymbol{\mathbf{x}})=\underset{\boldsymbol{\mathbf{u}}\in\mathcal{K}_{\alpha}(\boldsymbol{\mathbf{x}})}{\operatorname{argmin}}\>\|\boldsymbol{\mathbf{u}}-\kappa(\boldsymbol{\mathbf{x}})\|^{2},$ (12)
and $\kappa:\mathbb{R}^{n}\to\mathbb{R}^{m}$ is a nominal control policy with desirable closed-loop performance. When (9) holds, the optimization problem (12) is always feasible, making $\beta$ well-defined. If, in addition, the policy $\beta$ is locally Lipschitz continuous, there exists a unique solution $\boldsymbol{\mathbf{x}}(t)$ to the closed-loop system (11), and it satisfies $\boldsymbol{\mathbf{x}}(t)\!\in\!\mathcal{C}\!\subset\!\mathcal{X}$. In upcoming sections, we provide conditions under which $\beta$ is locally Lipschitz continuous.
Often, the sets $\mathcal{X}$ and $\mathcal{U}$ are the intersection of multiple state and input constraints. In these cases, it is usually easier to design a separate CBF for each constraint and then enforce all the conditions in one CBF-based program. Unfortunately, there is typically no guarantee that feasibility is retained when all the constraints are superimposed. The following property certifies feasibility is retained in these cases
Definition 2.4 ([9, Definition 2]).
Consider $n_{c}\in\mathbb{N}$ continuously differentiable functions $h_{i}:\mathcal{D}_{i}\to\mathbb{R}$ and let $\mathcal{C}_{i}=\{\boldsymbol{\mathbf{x}}\in\mathcal{D}_{i}\mid h_{i}(\boldsymbol{\mathbf{x}})\geq 0\}\subset\mathbb{R}^{n}$, with $\mathcal{C}=\cap_{i=1}^{n_{c}}\mathcal{C}_{i}$. The functions $h_{i}$ are said to have the control-sharing property if there exists a class $\mathscr{K}_{\infty}$ function $\alpha$ satisfying $\forall\boldsymbol{\mathbf{x}}\in\mathcal{C}$, $\exists\boldsymbol{\mathbf{u}}\in\mathcal{U}$, such that
$\min_{i\in\{1,\ldots,n_{c}\}}\left[L_{f}h_{i}(\boldsymbol{\mathbf{x}})+L_{g}h_{i}(\boldsymbol{\mathbf{x}})\boldsymbol{\mathbf{u}}+\alpha\big{(}h_{i}(\boldsymbol{\mathbf{x}})\big{)}\right]\!\geq\!0.$ (13)
In this paper, we call a smooth function $h:\mathcal{D}\to\mathbb{R}^{n_{c}}$ whose channels $h_{i}$ have the control-sharing property a (vector-valued) control barrier function.
The functions $h_{i}$ in Definition 2.4 are deliberately not labeled as CBFs individually (as opposed to [9, Definition 2]) because such a requirement is unnecessarily limiting. In fact, if 00 is a regular value of each $h_{i}$, Lemma 1 also holds for the set $\mathcal{C}$ defined as the intersection of the $\mathcal{C}_{i}$ even if each $\mathcal{C}_{i}$ is individually not compact (see [28, Proposition 4.3.7] for more details.) The optimization-based policy $\beta$ can also be used with (vector-valued) CBFs with the following admissible input set
[…]
As noted in the introduction, finding a CBF for arbitrary constraint sets $\mathcal{X}\subset\mathbb{R}^{n}$ is a challenging task. In fact, Lemma 1 states that, when $\mathcal{U}$ is compact, a CBF $h:\mathcal{D}\to\mathbb{R}$ satisfying $\mathcal{X}=\{\boldsymbol{\mathbf{x}}\in\mathcal{D}\mid h(\boldsymbol{\mathbf{x}})\geq 0\}$ with a well-defined policy $\beta$ exists only if $\mathcal{X}$ is control invariant, which is generally not the case. In spite of this, many practitioners insist on imposing $\mathcal{C}=\mathcal{X}$ to construct a “candidate” CBF (i.e. a continuously differentiable function $h:\mathcal{D}\to\mathbb{R}$ that satisfies $\mathcal{X}=\{\boldsymbol{\mathbf{x}}\in\mathcal{D}\mid h(\boldsymbol{\mathbf{x}})\geq 0\}$) and then tuning a class $\mathscr{K}_{\infty}$ function $\alpha$ on a case-by-case basis. Although this approach can yield good performance, it suffers from the possibility that $\mathcal{K}_{\alpha}(\boldsymbol{\mathbf{x}})=\emptyset$ for some $\boldsymbol{\mathbf{x}}\in\mathcal{X}$, which makes the controller $\beta$ ill-defined.
In this paper, we leverage results from the ERG framework to design valid CBFs that satisfy Definition 2 (and more generally Definition 2.4) and, additionally, are guaranteed to satisfy (9).
2.3 Dynamic Safety Margins
Dynamic safety margins are functions that quantify the distance to constraint violation for a prestabilized system subject to a constant reference. In order to use them, we must make the following assumptions about system (1).
Assumption 1
System (1) admits a simply connected equilibrium manifold parameterized by continuous functions $\bar{x}:\mathbb{R}^{l}\to\mathbb{R}^{n}$ and $\bar{u}:\mathbb{R}^{l}\to\mathbb{R}^{m}$ such that
$\forall\boldsymbol{\mathbf{v}}\in\mathbb{R}^{l},\quad f\big{(}\bar{x}(\boldsymbol{\mathbf{v}})\big{)}+g\big{(}\bar{x}(\boldsymbol{\mathbf{v}})\big{)}\bar{u}(\boldsymbol{\mathbf{v}})=0.$ (15)
Moreover, every equilibrium point $\bar{x}(\boldsymbol{\mathbf{v}})$ is stabilizable.
The vector $\boldsymbol{\mathbf{v}}\in\mathbb{R}^{l}$ is hereafter referred to as a “reference” for the system because it identifies a specific equilibrium point. The set of steady-state admissible references is
$\mathcal{V}=\{\boldsymbol{\mathbf{v}}\in\mathbb{R}^{l}\mid\bar{x}(\boldsymbol{\mathbf{v}})\in\mathcal{X},~{}\bar{u}(\boldsymbol{\mathbf{v}})\in\mathcal{U}\}.$ (16)
Note that, since $\bar{x}$, $\bar{u}$ are continuous and $\mathcal{X}$, $\mathcal{U}$ are closed sets, $\mathcal{V}$ is closed in $\mathbb{R}^{l}$.
Assumption 2
The set $\mathcal{V}\!\subset\!\mathbb{R}^{l}$ is not empty, i.e., system (1) features at least one steady-state admissible equilibrium point.
In essence, Assumption 1 establishes a continuous mapping from references to stabilizable equilibria, and Assumption 2 states that at least one of these equilibria is safe. By Assumption 1, let $\pi:\mathbb{R}^{n}\times\mathbb{R}^{l}\to\mathbb{R}^{m}$, locally Lipschitz continuous on $\mathbb{R}^{n}$ and continuous on $\mathbb{R}^{l}$, be a prestabilizing control law such that $\pi\big{(}\bar{x}(\boldsymbol{\mathbf{v}}),\boldsymbol{\mathbf{v}}\big{)}=\bar{u}(\boldsymbol{\mathbf{v}})$, and $\bar{x}(\boldsymbol{\mathbf{v}})$ is an asymptotically stable equilibrium point of the prestabilized system
$\dot{\boldsymbol{\mathbf{x}}}=f_{\pi}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\triangleq f(\boldsymbol{\mathbf{x}})+g(\boldsymbol{\mathbf{x}})\pi(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}}).$ (17)
Given a reference $\boldsymbol{\mathbf{v}}\in\mathbb{R}^{l}$, the prestabilized dynamics $f_{\pi}$ ensure the existence of an open set $\mathcal{D}_{\boldsymbol{\mathbf{v}}}\subset\mathbb{R}^{n}$ that satisfies $\bar{x}(\boldsymbol{\mathbf{v}})\in\mathcal{D}_{\boldsymbol{\mathbf{v}}}$ and is such that
$\boldsymbol{\mathbf{x}}(0)\in\mathcal{D}_{\boldsymbol{\mathbf{v}}}\quad\implies\quad\lim_{t\to\infty}\boldsymbol{\mathbf{x}}(t)=\bar{x}(\boldsymbol{\mathbf{v}}).$ (18)
The set $\mathcal{D}_{\boldsymbol{\mathbf{v}}}$ is a region of attraction for $\bar{x}(\boldsymbol{\mathbf{v}})$. Note that, for each reference $\boldsymbol{\mathbf{v}}\in\mathbb{R}^{l}$, the input constraint set $\mathcal{U}$ induces state constraints on the prestabilized system. This allows us to define the reference-dependent state constraint set
$\mathcal{X}_{\boldsymbol{\mathbf{v}}}=\{\boldsymbol{\mathbf{x}}\in\mathcal{X}\mid\pi(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\mathcal{U}\}.$ (19)
By continuity of $\pi$ on $\mathbb{R}^{l}$, it follows that $\mathcal{X}_{\boldsymbol{\mathbf{v}}}$ is closed in $\mathbb{R}^{n}$.
Let us now define the state-reference space $\mathbb{R}^{n}\times\mathbb{R}^{l}$ obtained by concatenating the state and reference vectors $[\boldsymbol{\mathbf{x}};~{}\boldsymbol{\mathbf{v}}]$.
Assumption 3
The state-reference region of attraction set
$\tilde{\mathcal{D}}\triangleq\{(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\mathbb{R}^{n}\times\mathbb{R}^{l}\mid\boldsymbol{\mathbf{x}}\in\mathcal{D}_{\boldsymbol{\mathbf{v}}}\},$ (20)
is open in $\mathbb{R}^{n}\times\mathbb{R}^{l}$.
We also define the state-reference constraint set
$\tilde{\mathcal{X}}\triangleq\{(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\mathbb{R}^{n}\times\mathbb{R}^{l}\mid\boldsymbol{\mathbf{x}}\in\mathcal{X}_{\boldsymbol{\mathbf{v}}},\boldsymbol{\mathbf{v}}\in\mathcal{V}\}.$ (21)
The following DSM definition satisfies the one proposed in [23, Definition 1], although it makes stronger assumptions by requiring a) differentiability rather than continuity and b) invariance rather than strong returnability.
Definition 2.5.
Let $\Delta:\tilde{\mathcal{D}}\to\mathbb{R}^{n_{c}}$ be continuously differentiable and define the set $\tilde{\mathcal{C}}=\{(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\tilde{\mathcal{D}}\mid\Delta(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\geq 0\}$. The function $\Delta$ is called a dynamic safety margin (DSM) for $\pi$ if the following conditions hold
$\tilde{\mathcal{C}}\subset\tilde{\mathcal{X}},$ (22a)
[…]
$\Delta_{i}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})=0~{}\implies~{}\frac{\partial\Delta_{i}}{\partial\boldsymbol{\mathbf{x}}}f_{\pi}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\geq 0,$ (22c)
where the last implication holds for all $i\in\{1,\ldots,n_{c}\}$.
DSMs are used to design add-on modules (ERGs) that enforce constraint satisfaction by filtering the reference of the prestabilized system. Specifically, let us augment the prestabilized system with reference dynamics to obtain
$\begin{bmatrix}\dot{\boldsymbol{\mathbf{x}}}\\ \dot{\boldsymbol{\mathbf{v}}}\end{bmatrix}=\begin{bmatrix}f(\boldsymbol{\mathbf{x}})+g(\boldsymbol{\mathbf{x}})\pi(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\\ \min_{i\in\{1,\ldots,n_{c}\}}\Delta_{i}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\rho(\boldsymbol{\mathbf{v}},\boldsymbol{\mathbf{r}})\end{bmatrix},$ (23)
where $\rho:\mathbb{R}^{l}\times\mathbb{R}^{l}\to\mathbb{R}^{l}$ is any locally Lipschitz continuous function such that $\boldsymbol{\mathbf{r}}\in\mathbb{R}^{l}$ is an asymptotically stable equilibrium point of the system $\dot{\boldsymbol{\mathbf{v}}}(t)=\rho\big{(}\boldsymbol{\mathbf{v}}(t),\boldsymbol{\mathbf{r}}\big{)}$. In the ERG literature, $\rho$ is called a navigation field [23, Definition 2]. System (23) ensures that, if $\big{(}\boldsymbol{\mathbf{x}}(0),\boldsymbol{\mathbf{v}}(0)\big{)}\in\tilde{\mathcal{C}}$, then $\big{(}\boldsymbol{\mathbf{x}}(t),\boldsymbol{\mathbf{v}}(t)\big{)}\in\tilde{\mathcal{C}}\subset\tilde{\mathcal{X}}$ at all times [23]. The ERG literature provides systematic tools for computing DSMs for various classes of systems [30, 23, 31, 32, 33, 34]. The following section shows how these tools from the ERG literature can be used to construct valid CBFs.
3 From Dynamic Safety Margins to Control Barrier Functions
This section explores the connections between DSMs and CBFs, formulates a safe, optimization-based control policy and provides conditions for this policy to be locally Lipschitz continuous.
3.1 DSMs are CBFs
Based on the intuition that both CBFs and DSMs can be used to design add-on modules for constraint handling, the following theorem states that, a dynamic safety margin is, in fact, a CBF for an augmented system consisting of the concatenation of $\boldsymbol{\mathbf{x}}\in\mathbb{R}^{n}$ and $\boldsymbol{\mathbf{v}}\in\mathbb{R}^{l}$
$\begin{bmatrix}\dot{\boldsymbol{\mathbf{x}}}\\ \dot{\boldsymbol{\mathbf{v}}}\end{bmatrix}=\begin{bmatrix}f(\boldsymbol{\mathbf{x}})+g(\boldsymbol{\mathbf{x}})\boldsymbol{\mathbf{u}}\\ \boldsymbol{\mathbf{w}}\end{bmatrix},$ (24)
where $(\boldsymbol{\mathbf{u}},\boldsymbol{\mathbf{w}})\in\mathbb{R}^{m}\times\mathbb{R}^{l}$ are the augmented inputs.
Theorem 3.6 (DSMs are CBFs).
If $\Delta:\tilde{\mathcal{D}}\to\mathbb{R}^{n_{c}}$ is a DSM for $\pi$, then $\Delta$ is a (vector-valued) CBF with respect to the augmented system (24).
Proof 3.7.
Define the set-valued maps
$\tilde{\mathcal{C}}_{i}(c)=\{(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\tilde{\mathcal{C}}:0\leq\Delta_{i}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\leq c\},$ (25)
and the functions $\hat{\alpha}_{i}:[0,\infty)\to\mathbb{R}$ by
$\hat{\alpha}_{i}(c)=-\inf_{(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\tilde{\mathcal{C}}_{i}(c)}\frac{\partial\Delta_{i}}{\partial\boldsymbol{\mathbf{x}}}f_{\pi}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}}),$ (26)
where $i\in\{1,\ldots,n_{c}\}$. Note that for any $c\in[0,\infty)$, the set $\tilde{\mathcal{C}}_{i}(c)\subset\tilde{\mathcal{C}}$ is compact. It can also be shown that the set-valued maps $\mathcal{C}_{i}$ are both upper- and lower-hemicontinuous. Thus, it follows by Berge’s maximum theorem [29] that each $\hat{\alpha}_{i}$ is continuous on $[0,\infty)$. Moreover, it is nondecreasing because for any $c_{1}\leq c_{2}$, $\tilde{\mathcal{C}}_{i}(c_{1})\subset\tilde{\mathcal{C}}_{i}(c_{2})$. Next, note that for any $(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\tilde{\mathcal{C}}_{i}(0)$, $\Delta_{i}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})=0$, and it follows by (22c) that
$\frac{\partial\Delta_{i}}{\partial\boldsymbol{\mathbf{x}}}f_{\pi}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\geq 0.$
Therefore,
$\hat{\alpha}_{i}(0)=-\inf_{(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\tilde{\mathcal{C}}_{i}(0)}\frac{\partial\Delta_{i}}{\partial\boldsymbol{\mathbf{x}}}f_{\pi}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\leq 0.$ (27)
There always exists $\alpha_{i}\in\mathscr{K}_{\infty}$ such that $\forall c\in[0,\infty)$, $\hat{\alpha}_{i}(c)\leq\alpha_{i}(c)$. Define $\alpha(c)=\max_{i\in\{1,\ldots,n_{c}\}}\alpha_{i}(c)$ and note that $\alpha\in\mathscr{K}_{\infty}$. Let $(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\tilde{\mathcal{C}}$ and $i\in\{1,\ldots,n_{c}\}$ be given and note that since $\boldsymbol{\mathbf{x}}\in\mathcal{X}_{\boldsymbol{\mathbf{v}}}$, it follows that $\pi(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\mathcal{U}$. With this,
$\dot{\Delta}_{i}\big{(}\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}},\pi(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}}),0\big{)}$ $=\frac{\partial\Delta_{i}}{\partial\boldsymbol{\mathbf{x}}}f_{\pi}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})+\frac{\partial\Delta_{i}}{\partial\boldsymbol{\mathbf{v}}}0$
$\geq\inf_{(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\tilde{\mathcal{C}}_{i}(\Delta_{i}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}}))}\frac{\partial\Delta_{i}}{\partial\boldsymbol{\mathbf{x}}}f_{\pi}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})$
$\geq-\hat{\alpha}_{i}\big{(}\Delta_{i}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\big{)}$
$\geq-\alpha_{i}\big{(}\Delta_{i}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\big{)}$
$\geq-\alpha\big{(}\Delta_{i}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\big{)}.$
We have shown that $\forall(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\tilde{\mathcal{C}}$, $(\boldsymbol{\mathbf{u}},\boldsymbol{\mathbf{w}})=\big{(}\pi(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}}),0\big{)}\in\mathcal{U}\times\mathbb{R}^{l}$ is such that
$\min_{i\in\{1,\ldots,n_{c}\}}\left[\dot{\Delta}_{i}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}},\boldsymbol{\mathbf{u}},\boldsymbol{\mathbf{w}})+\alpha\big{(}\Delta_{i}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\big{)}\right]\geq 0.$ (28)
We conclude that the channels $\Delta_{i}$ have the control-sharing property, so $\Delta$ is a (vector-valued) CBF.
Let us define the set of viable inputs at $(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})$ for DSM-CBFs as
[…]
The following corollary states that DSM-CBFs satisfy (9), so $\mathcal{U}$ needs not be compact to conclude control invariance of $\tilde{\mathcal{C}}$.
Corollary 3.8.
Let $\Delta:\tilde{\mathcal{D}}\to\mathbb{R}^{n_{c}}$ be a DSM for $\pi$. Then, there exists $\alpha\in\mathscr{K}_{\infty}$ such that $\Delta$ satisfies (9). Specifically, $\forall(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\tilde{\mathcal{C}}$, $\big{(}\pi(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}}),0\big{)}\in\tilde{\mathcal{K}}_{\alpha}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})$. Moreover, if 00 is a regular value of each $\Delta_{i}$, then the set $\tilde{\mathcal{C}}\subset\tilde{\mathcal{X}}\subset\mathcal{X}\times\mathcal{V}$ is control invariant.
Proof 3.9.
The first statement was already shown in the proof of Theorem 3.6. The control invariance of $\tilde{\mathcal{C}}$ is an immediate consequence of Theorem 2 when $n_{c}=1$. For the general case $n_{c}>1$, we refer the reader to [28, Proposition 4.3.7].
A significant concern in the CBF literature arises when the term $L_{g}h(\boldsymbol{\mathbf{x}})=0$ for some $\boldsymbol{\mathbf{x}}\in\mathcal{C}$. This means $h$ lacks uniform relative degree 1 over $\mathcal{C}$, and implies a loss of controllability that can be catastrophic when $L_{f}h(\boldsymbol{\mathbf{x}})<-\alpha\big{(}h(\boldsymbol{\mathbf{x}})\big{)}$. Although many relevant works address this issue (e.g. [7, 8]), and Corollary 3.8 already shows it is not a concern for DSM-CBFs, the following corollary shows an additional desirable property of DSM-CBFs under total controllability loss.
Corollary 3.10.
Let $\Delta:\tilde{\mathcal{D}}\to\mathbb{R}^{n_{c}}$ be a DSM for $\pi$ and let $\alpha\in\mathscr{K}_{\infty}$ be such that (9) holds. If $(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\tilde{\mathcal{C}}$ is such that
$\begin{bmatrix}\frac{\partial\Delta}{\partial\boldsymbol{\mathbf{x}}}g(\boldsymbol{\mathbf{x}})&\frac{\partial\Delta}{\partial\boldsymbol{\mathbf{v}}}\end{bmatrix}=0,$ (30)
then, $\tilde{\mathcal{K}}_{\alpha}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})=\mathcal{U}\times\mathbb{R}^{l}$.
Proof 3.11.
Note that since $\big{(}\pi(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}}),0\big{)}\in\tilde{\mathcal{K}}_{\alpha}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})$, it follows that $\forall i\in\{1,\ldots,n_{c}\}$,
$\dot{\Delta}_{i}\big{(}\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}},\pi(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}}),0\big{)}\!$ $=\!\frac{\partial\Delta_{i}}{\partial\boldsymbol{\mathbf{x}}}f(\boldsymbol{\mathbf{x}})\!+\!\begin{bmatrix}\frac{\partial\Delta_{i}}{\partial\boldsymbol{\mathbf{x}}}g(\boldsymbol{\mathbf{x}})\!&\!\frac{\partial\Delta_{i}}{\partial\boldsymbol{\mathbf{v}}}\end{bmatrix}\!\begin{bmatrix}\pi(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\\ 0\end{bmatrix}$
$=\frac{\partial\Delta_{i}}{\partial\boldsymbol{\mathbf{x}}}f(\boldsymbol{\mathbf{x}})\geq-\alpha\big{(}\Delta_{i}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\big{)}.$
Therefore, constraint (28) is trivially satisfied at $(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})$ for any augmented input $(\boldsymbol{\mathbf{u}},\boldsymbol{\mathbf{w}})$. We conclude that $\tilde{\mathcal{K}}_{\alpha}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})=\mathcal{U}\times\mathbb{R}^{l}$.
3.2 DSM-CBF Optimization-based Policy
Recalling Theorem 3.6, it is possible to reformulate the CBF policy $\beta$ for the augmented system $(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})$. To do so, let $\boldsymbol{\mathbf{r}}\in\mathbb{R}^{l}$ be a desired reference, let $\kappa:\mathbb{R}^{n}\times\mathbb{R}^{l}\to\mathbb{R}^{m}$ be a nominal control policy with desirable closed-loop performance, and let $\rho:\mathbb{R}^{l}\times\mathbb{R}^{l}\to\mathbb{R}^{l}$ be a navigation field. The following theorem formulates the CBF-based policy $\beta$ for DSMs and the augmented system (24).
Theorem 3.12 (DSM-CBF policy).
Let $\eta>0$ be a constant scalar. If $\Delta:\tilde{\mathcal{D}}\to\mathbb{R}^{n_{c}}$ is a DSM, then there exists $\alpha\in\mathscr{K}_{\infty}$ such that the optimization-based policy $\beta:\tilde{\mathcal{C}}\to\mathcal{U}\times\mathbb{R}^{l}$
$\beta(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})=\!\underset{(\boldsymbol{\mathbf{u}},\boldsymbol{\mathbf{w}})\in\tilde{\mathcal{K}}_{\alpha}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})}{\operatorname{argmin}}\|\boldsymbol{\mathbf{u}}\!-\!\kappa(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{r}})\|^{2}\!+\eta\|\boldsymbol{\mathbf{w}}\!-\!\rho(\boldsymbol{\mathbf{v}},\boldsymbol{\mathbf{r}})\|^{2},$ (31)
is feasible for all $(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\tilde{\mathcal{C}}$. Moreover, if $\beta$ is locally Lipschitz continuous and 00 is a regular value of each $\Delta_{i}$, then the trajectory of the augmented system (24) under policy $(\boldsymbol{\mathbf{u}},\boldsymbol{\mathbf{w}})=\beta(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})$ exists, is unique, and satisfies $\boldsymbol{\mathbf{x}}(t)\in\mathcal{X}$ and $\boldsymbol{\mathbf{v}}(t)\in\mathcal{V}$ for all $t\geq 0$, given $\big{(}\boldsymbol{\mathbf{x}}(0),\boldsymbol{\mathbf{v}}(0)\big{)}\in\tilde{\mathcal{C}}$.
Proof 3.13.
Feasibility of the DSM-CBF policy follows by Corollary 3.8. Any locally Lipschitz controller in $\tilde{\mathcal{K}}_{\alpha}$ certifies the local (in time) existence and uniqueness of the solution to the closed-loop system [35]. Due to the the invariance of $\tilde{\mathcal{C}}$, there can be no finite-time escape instances and the existence and uniqueness holds for all $t\geq 0$. Safety follows from $\tilde{\mathcal{C}}\subset\mathcal{X}\times\mathcal{V}$.
Remark 3.14.
Convexity of the set $\tilde{\mathcal{K}}_{\alpha}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})$, and therefore problem (31), is linked to convexity of $\mathcal{U}$. In particular, if the input constraint set $\mathcal{U}$ is polyhedral, the optimization problem (31) becomes a quadratic program (QP) for all $(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\tilde{\mathcal{C}}$, which can be solved efficiently.
Remark 3.15.
The positive constant $\eta>0$ makes the objective function of the augmented CBF-program (31) strictly convex. Moreover, we observed in practice that as $\eta\to\infty$, the system behavior approaches that of the ERG, which tends to have worse performance than can be achieved with $\eta\to 0$.
Remark 3.16.
If $\mathcal{V}$ is convex, a suitable navigation field is $\rho(\boldsymbol{\mathbf{v}},\boldsymbol{\mathbf{r}})=\boldsymbol{\mathbf{r}}-\boldsymbol{\mathbf{v}}$. Otherwise, please refer to works like [36].
Remark 3.17.
The proposed approach can be interpreted as a closed-form variant of the backup CBF [21]. The approach assigns a safe control input $\boldsymbol{\mathbf{u}}$ while simultaneously generating a safe reference $\boldsymbol{\mathbf{v}}$ for the backup controller $\pi$.
Remark 3.18.
Note that the nominal controller $\kappa$ and the prestabilizing controller $\pi$ can be different. In practice, we observed that mild prestabilizing controllers $\pi$ lead to larger feasible regions on the CBF-based program, allowing the safe input to be closer to the nominal target $\kappa$.
3.3 Existence and Uniqueness of Solutions
Given the prevalence of optimization-based policies in constrained control approaches such as MPC and CBFs, numerous works have studied their regularity properties [37, 12, 38].
Given any initial condition $(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\tilde{\mathcal{C}},$ the optimization-based policy $\beta$ for DSM-based CBFs, defined in (31), is always feasible. If, additionally, it is locally Lipschitz continuous, then the trajectory for the closed-loop system
$\begin{bmatrix}\dot{\boldsymbol{\mathbf{x}}}\\ \dot{\boldsymbol{\mathbf{v}}}\end{bmatrix}=\begin{bmatrix}f(\boldsymbol{\mathbf{x}})\\ 0\end{bmatrix}+\begin{bmatrix}g(\boldsymbol{\mathbf{x}})&0\\ 0&I\end{bmatrix}\beta(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}}),$ (32)
exists and is unique. Based on [38], we can formulate a reduced set of sufficient conditions that guarantee $\beta$ is locally Lipschitz continuous.
Theorem 3.19 ([38, Theorem 5.2]).
Let $\Delta:\tilde{\mathcal{D}}\to\mathbb{R}^{n_{c}}$ be a DSM for $\pi$, let $\rho$ be a navigation field and let $\eta>0$. Assume the input constraints are polyhedral $\mathcal{U}=\{\boldsymbol{\mathbf{u}}\in\mathbb{R}^{m}:M\boldsymbol{\mathbf{u}}\leq\boldsymbol{\mathbf{b}}\}$, and the functions $\pi(\cdot,\boldsymbol{\mathbf{r}})$ and $\rho(\cdot,\boldsymbol{\mathbf{r}})$ are continuously differentiable. If these constraint qualifications (CQs) hold
  1. a)
    (Slater CQ) There exists $\alpha\in\mathscr{K}_{\infty}$ such that for all $(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\tilde{\mathcal{C}}$, $\operatorname{Int}\left(\tilde{\mathcal{K}}_{\alpha}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\right)\neq\emptyset$;
  2. b)
    (Constant Rank CQ) There exists a neighborhood of every point in $\tilde{\mathcal{C}}$ wherein, for any subset $\mathcal{I}\subset\{1,\ldots,n_{c}\}$, the number of linearly independent vectors in $\left\{\begin{bmatrix}\frac{\partial\Delta_{i}}{\partial\boldsymbol{\mathbf{x}}}g(\boldsymbol{\mathbf{x}})&\frac{\partial\Delta_{i}}{\partial\boldsymbol{\mathbf{v}}}\end{bmatrix},i\in\mathcal{I}\right\}$ remains constant;
then the optimization-based policy $\beta$ is locally Lipschitz continuous and the solution to the closed-loop system (32) exists and is unique.
Remark 3.20.
In the special case that $n_{c}=1$, both Slater and Constant Rank CQs hold if $\Delta$ has uniform relative degree $1$ over $\tilde{\mathcal{C}}$ and $\pi(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\operatorname{Int}\left(\mathcal{U}\right)$ for all $(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\tilde{\mathcal{C}}$.
4 Lyapunov-based Dynamic Safety Margins
This section shows that the Lyapunov-based DSMs in [23, 30, 31, 32] can be used within the context of this paper.
Definition 4.21.
A continuously differentiable function $V:\mathbb{R}^{n}\times\mathbb{R}^{l}\to\mathbb{R}$ is a reference-dependent Lyapunov function for the prestabilized system $f_{\pi}$ if there exist functions $\gamma_{1},\gamma_{2}\in\mathscr{K}_{\infty}$ and $\forall\boldsymbol{\mathbf{v}}\in\mathbb{R}^{l}$, there exists a neighborhood of $\bar{x}(\boldsymbol{\mathbf{v}})$, denoted $\mathcal{D}_{\boldsymbol{\mathbf{v}}}\subset\mathbb{R}^{n}$, wherein
$\gamma_{1}\big{(}\|\boldsymbol{\mathbf{x}}-\bar{x}(\boldsymbol{\mathbf{v}})\|\big{)}$ $\leq V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\leq\gamma_{2}\big{(}\|\boldsymbol{\mathbf{x}}-\bar{x}(\boldsymbol{\mathbf{v}})\|\big{)},$ (33a)
$\frac{\partial V}{\partial\boldsymbol{\mathbf{x}}}f_{\pi}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})$ $\leq 0.$ (33b)
Remark 4.22.
Note that the reference-dependent Lyapunov function $V$ needs not certify asymptotic stability of the equilibria of the prestabilized system $f_{\pi}$. That is, $\dot{V}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})=0$ is acceptable for $\boldsymbol{\mathbf{x}}\neq\bar{x}(\boldsymbol{\mathbf{v}})$. This allows for more general classes of Lyapunov functions to be considered and ultimately simplifies the design of the DSM.
Additionally, we require $V$ has non-vanishing gradient near the equilibrium points. Formally, we have the following assumption.
Assumption 4
Given any $\boldsymbol{\mathbf{v}}\in\mathbb{R}^{l}$, $\boldsymbol{\mathbf{x}}\in\mathcal{D}_{\boldsymbol{\mathbf{v}}}\setminus\{\bar{x}(\boldsymbol{\mathbf{v}})\}$ is a regular point of $V(\cdot,\boldsymbol{\mathbf{v}}):\mathbb{R}^{n}\to\mathbb{R}$.
Recalling the reference-dependent state constraint set $\mathcal{X}_{\boldsymbol{\mathbf{v}}}$ defined in (19), let $\mathcal{X}^{c}_{\boldsymbol{\mathbf{v}}}=\mathbb{R}^{n}\setminus\mathcal{X}_{\boldsymbol{\mathbf{v}}}$ be its complement, which is open in $\mathbb{R}^{n}$. We define the safety threshold value $\Gamma^{*}:\mathbb{R}^{l}\to\mathbb{R}$
$\Gamma^{*}(\boldsymbol{\mathbf{v}})\triangleq\inf_{\boldsymbol{\mathbf{x}}\in\>\mathcal{X}^{c}_{\boldsymbol{\mathbf{v}}}\cap\mathcal{D}_{\boldsymbol{\mathbf{v}}}}V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})-\inf_{\boldsymbol{\mathbf{x}}\in\mathcal{X}_{\boldsymbol{\mathbf{v}}}\cap\mathcal{D}_{\boldsymbol{\mathbf{v}}}}V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}}).$ (34)
When positive, $\Gamma^{*}(\boldsymbol{\mathbf{v}})$ represents the largest constraint-admissible level set of the Lyapunov function $V$ at reference $\boldsymbol{\mathbf{v}}$. Our definition for $\Gamma^{*}$ is an extension of the threshold value in [23] to the entire reference space $\mathbb{R}^{l}$. With this, $\Gamma^{*}$ has the following property.
Lemma 4.23.
Let $V:\mathbb{R}^{n}\times\mathbb{R}^{l}\to\mathbb{R}$ be a reference-dependent Lyapunov function that satisfies Assumption 4. We have that $\Gamma^{*}(\boldsymbol{\mathbf{v}})\geq 0$ if and only if $\boldsymbol{\mathbf{v}}\in\mathcal{V}$.
Proof 4.24.
Let $\boldsymbol{\mathbf{v}}\in\mathcal{V}$ be given. Then, $\bar{x}(\boldsymbol{\mathbf{v}})\in\mathcal{X}_{\boldsymbol{\mathbf{v}}}\cap\mathcal{D}_{\boldsymbol{\mathbf{v}}}$, which implies that $\inf_{\boldsymbol{\mathbf{x}}\in\mathcal{X}_{\boldsymbol{\mathbf{v}}}\cap\mathcal{D}_{\boldsymbol{\mathbf{v}}}}V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})=0$. Recalling that $\forall\boldsymbol{\mathbf{x}}\in\mathcal{D}_{\boldsymbol{\mathbf{v}}}$, $V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\geq 0$, we conclude that $\Gamma^{*}(\boldsymbol{\mathbf{v}})\geq 0$.
For the other direction, we show the contrapositive statement. Let $\boldsymbol{\mathbf{v}}\in\mathbb{R}^{l}\setminus\mathcal{V}$ be given. Then, $\bar{x}(\boldsymbol{\mathbf{v}})\in\mathcal{X}_{\boldsymbol{\mathbf{v}}}^{c}\cap\mathcal{D}_{\boldsymbol{\mathbf{v}}}$, which implies that $\inf_{\boldsymbol{\mathbf{x}}\in\mathcal{X}_{\boldsymbol{\mathbf{v}}}^{c}\cap\mathcal{D}_{\boldsymbol{\mathbf{v}}}}V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})=0$. Since $\mathcal{X}_{\boldsymbol{\mathbf{v}}}^{c}\cap\mathcal{D}_{\boldsymbol{\mathbf{v}}}$ is open in $\mathbb{R}^{n}$, there exists $\delta>0$ such that $\mathcal{B}\big{(}\bar{x}(\boldsymbol{\mathbf{v}}),\delta\big{)}=\{\boldsymbol{\mathbf{x}}\in\mathbb{R}^{n}\mid\|\boldsymbol{\mathbf{x}}-\bar{x}(\boldsymbol{\mathbf{v}})\|<\delta\}\subset\mathcal{X}_{\boldsymbol{\mathbf{v}}}^{c}\cap\mathcal{D}_{\boldsymbol{\mathbf{v}}}$. Thus,
$\Gamma^{*}(\boldsymbol{\mathbf{v}})=-\inf_{\boldsymbol{\mathbf{x}}\in\mathcal{X}_{\boldsymbol{\mathbf{v}}}\cap\mathcal{D}_{\boldsymbol{\mathbf{v}}}}V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\leq-\gamma_{1}(\delta)<0.$ (35)
Let us also define the stability threshold value $\overline{\Gamma}:\mathbb{R}^{l}\to\mathbb{R}$
$\overline{\Gamma}(\boldsymbol{\mathbf{v}})\triangleq\inf_{\boldsymbol{\mathbf{x}}\in\partial\mathcal{D}_{\boldsymbol{\mathbf{v}}}}V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}}).$ (36)
We can now construct DSMs using Lyapunov functions.
Theorem 4.25 (Lyapunov-based DSMs).
Let $V:\mathbb{R}^{n}\times\mathbb{R}^{l}\to\mathbb{R}$ be a reference-dependent Lyapunov function for the prestabilized system $f_{\pi}$ satisfying Assumption 4. If $\Gamma^{*}:\mathbb{R}^{l}\to\mathbb{R}$ and $\overline{\Gamma}:\mathbb{R}^{l}\to\mathbb{R}$ are continuously differentiable, and $\mathcal{X}\subset\mathbb{R}^{n}$ and $\mathcal{V}\subset\mathbb{R}^{l}$ are compact, then for any $\epsilon\in(0,1)$,
$\Delta(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})=\begin{bmatrix}\Gamma^{*}(\boldsymbol{\mathbf{v}})-V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\\ (1-\epsilon)\overline{\Gamma}(\boldsymbol{\mathbf{v}})-V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\end{bmatrix},$ (37)
is a dynamic safety margin for $\pi$.
Proof 4.26.
To prove property (22a), let $(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\tilde{\mathcal{C}}\subset\tilde{\mathcal{D}}$. Then, $\Gamma^{*}(\boldsymbol{\mathbf{v}})\geq V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\geq 0$, which implies that $\boldsymbol{\mathbf{v}}\in\mathcal{V}$ by Lemma 4.23. Next, for a contradiction, assume $\boldsymbol{\mathbf{x}}\notin\mathcal{X}_{\boldsymbol{\mathbf{v}}}$. Then
$\Gamma^{*}(\boldsymbol{\mathbf{v}})=\inf_{\boldsymbol{\mathbf{x}}\in\mathcal{X}_{\boldsymbol{\mathbf{v}}}^{c}\cap\mathcal{D}_{\boldsymbol{\mathbf{v}}}}V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\leq V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}}),$
which implies that $\Gamma^{*}(\boldsymbol{\mathbf{v}})\leq\Gamma(\boldsymbol{\mathbf{v}})\leq V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})$. So, we must have $V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})=\Gamma^{*}(\boldsymbol{\mathbf{v}})$. If $V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})=0$, then $\boldsymbol{\mathbf{x}}=\bar{x}(\boldsymbol{\mathbf{v}})\in\mathcal{X}_{\boldsymbol{\mathbf{v}}}$ because $\boldsymbol{\mathbf{v}}\in\mathcal{V}$, and we have a contradiction. Otherwise, if $\Gamma^{*}(\boldsymbol{\mathbf{v}})=V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})>0$, recall that $\mathcal{X}^{c}_{\boldsymbol{\mathbf{v}}}\cap\mathcal{D}_{\boldsymbol{\mathbf{v}}}$ is open in $\mathbb{R}^{n}$. Therefore, there exists an open set $\mathcal{N}\subset\mathcal{X}^{c}_{\boldsymbol{\mathbf{v}}}\cap\mathcal{D}_{\boldsymbol{\mathbf{v}}}$ with $\boldsymbol{\mathbf{x}}\in\mathcal{N}$. Furthermore, since $\boldsymbol{\mathbf{x}}\in\mathcal{D}_{\boldsymbol{\mathbf{v}}}\setminus\{\bar{x}(\boldsymbol{\mathbf{v}})\}$ is a regular point of $V(\cdot,\boldsymbol{\mathbf{v}})$ by Assumption 4, we have that $\frac{\partial V}{\partial\boldsymbol{\mathbf{x}}}\neq 0$. So, there exists a sufficiently small scalar $\delta>0$ such that $\hat{\boldsymbol{\mathbf{x}}}=\boldsymbol{\mathbf{x}}-\delta\frac{\partial V}{\partial\boldsymbol{\mathbf{x}}}^{\top}\in\mathcal{N}\subset\mathcal{X}^{c}_{\boldsymbol{\mathbf{v}}}\cap\mathcal{D}_{\boldsymbol{\mathbf{v}}}$. Furthermore, this point achieves lower value in $V$. That is, $V(\hat{\boldsymbol{\mathbf{x}}},\boldsymbol{\mathbf{v}})<V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})$. So, $\Gamma^{*}(\boldsymbol{\mathbf{v}})=V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})$ is a contradiction because
$\Gamma^{*}(\boldsymbol{\mathbf{v}})=\inf_{\boldsymbol{\mathbf{x}}\in\mathcal{X}^{c}_{\boldsymbol{\mathbf{v}}}\cap\mathcal{D}_{\boldsymbol{\mathbf{v}}}}V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\leq V(\hat{\boldsymbol{\mathbf{x}}},\boldsymbol{\mathbf{v}})<V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}}).$
To prove property (22b), note that since $\tilde{\mathcal{C}}\subset\mathcal{X}\times\mathcal{V}$, and $\mathcal{X}\times\mathcal{V}$ is compact in $\mathbb{R}^{n}\times\mathbb{R}^{l}$, we must only show that $\tilde{\mathcal{C}}$ is closed in $\mathbb{R}^{n}\times\mathbb{R}^{l}$. Let $\{(\boldsymbol{\mathbf{x}}_{n},\boldsymbol{\mathbf{v}}_{n})\}\subset\tilde{\mathcal{C}}$ be a sequence that converges to $(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\mathbb{R}^{n}\times\mathbb{R}^{l}$. Since $\mathcal{X}\subset\mathbb{R}^{n}$ and $\mathcal{V}\subset\mathbb{R}^{l}$ are closed, it follows that $\boldsymbol{\mathbf{x}}\in\mathcal{X}$ and $\boldsymbol{\mathbf{v}}\in\mathcal{V}$. Now we argue $\Delta(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\geq 0$. For a contradiction, assume there is $i\in\{1,2\}$ such that $\Delta_{i}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})<0$. By continuity of $\Delta_{i}$ on $\mathbb{R}^{n}\times\mathbb{R}^{l}$, there exists an open neighborhood $\tilde{\mathcal{N}}\subset\mathbb{R}^{n}\times\mathbb{R}^{l}$ of $(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})$ wherein $\Delta_{i}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})<0$. By convergence of $\{(\boldsymbol{\mathbf{x}}_{n},\boldsymbol{\mathbf{v}}_{n})\}$ to $(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})$, there exists $N\in\mathbb{N}$ such that $(\boldsymbol{\mathbf{x}}_{N},\boldsymbol{\mathbf{v}}_{N})\in\tilde{\mathcal{N}}$, which raises a contradiction because $(\boldsymbol{\mathbf{x}}_{N},\boldsymbol{\mathbf{v}}_{N})\in\tilde{\mathcal{C}}$ implies that $\Delta(\boldsymbol{\mathbf{x}}_{N},\boldsymbol{\mathbf{v}}_{N})\geq 0$. It remains to show that $(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\tilde{\mathcal{D}}$. Begin by noting that for all $n\in\mathbb{N}$, $(\boldsymbol{\mathbf{x}}_{n},\boldsymbol{\mathbf{v}}_{n})\in\tilde{\mathcal{D}}$ implies that $V(\boldsymbol{\mathbf{x}}_{n},\boldsymbol{\mathbf{v}}_{n})\geq\gamma_{1}\big{(}\|\boldsymbol{\mathbf{x}}_{n}-\bar{x}(\boldsymbol{\mathbf{v}}_{n})\|\big{)}$. It follows by continuity of all functions involved in the inequality and by the order limit theorem that $V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\geq\gamma_{1}\big{(}\|\boldsymbol{\mathbf{x}}-\bar{x}(\boldsymbol{\mathbf{v}})\|\big{)}$. For a contradiction, let us assume that $(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\partial\tilde{\mathcal{D}}$, which implies that $\boldsymbol{\mathbf{x}}\in\partial\mathcal{D}_{\boldsymbol{\mathbf{v}}}$ and, therefore, $\overline{\Gamma}(\boldsymbol{\mathbf{v}})\leq V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})$. However, recall that we have shown $\Delta(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\geq 0$. Thus, $V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\leq(1-\epsilon)\overline{\Gamma}(\boldsymbol{\mathbf{v}})$. This yields
$\frac{1}{1-\epsilon}V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\leq\bar{\Gamma}(\boldsymbol{\mathbf{v}})\leq V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}}),$
which is a contradiction because $V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\geq\gamma_{1}\big{(}\|\boldsymbol{\mathbf{x}}-\bar{x}(\boldsymbol{\mathbf{v}})\|\big{)}>0$. So, we must have that $(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\tilde{\mathcal{D}}$ and we conclude that $\tilde{\mathcal{C}}$ is closed in $\mathbb{R}^{n}\times\mathbb{R}^{l}$, and therefore compact.
To prove property (22c), let $i\in\{1,2\}$ be given and pick $(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\tilde{\mathcal{D}}$ such that $\Delta_{i}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})=0$. Then, by (33b),
$\frac{\partial\Delta_{i}}{\partial\boldsymbol{\mathbf{x}}}f_{\pi}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})=-\frac{\partial V}{\partial\boldsymbol{\mathbf{x}}}f_{\pi}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\geq 0.$ (38)
The following corollary highlights that for Lyapunov-based DSMs, any $\alpha\in\mathscr{K}_{\infty}$ can be used for the DSM-CBF policy $\beta$ given in (31). This result constitutes a step away from existence-type proofs common in the CBF literature and towards implementation.
Corollary 4.27.
If $\Delta$ is a Lyapunov-based DSM as in (37), then $\forall\alpha\in\mathscr{K}_{\infty}$, $\forall(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\tilde{\mathcal{C}}$, $\tilde{\mathcal{K}}_{\alpha}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\neq\emptyset$.
Proof 4.28.
Let $\alpha\in\mathscr{K}_{\infty}$ and $(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\tilde{\mathcal{C}}$ be given. Recall that $\pi(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\mathcal{U}$ and note that for any $i\in\{1,2\}$,
$\dot{\Delta}_{i}\big{(}\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}},\pi(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}}),0\big{)}=-\frac{\partial V}{\partial\boldsymbol{\mathbf{x}}}f_{\pi}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\geq 0\geq-\alpha\big{(}\Delta_{i}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\big{)}.$
Therefore, $\big{(}\pi(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}}),0\big{)}\in\tilde{\mathcal{K}}_{\alpha}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})$.
Remark 4.29.
When multiple constraints are considered, it is usually simpler to define multiple safety threshold values $\Gamma^{*}_{i}$ and stack them into the form of (37). As long as each one is differentiable, Theorem 4.25 can still be applied to obtain a DSM $\Delta:\tilde{\mathcal{D}}\to\mathbb{R}^{n_{c}}$ with $n_{c}>2$. In the event either $\Gamma^{*}_{i}$ or $\overline{\Gamma}$ are not continuously differentiable, any continuously differentiable, lower-bounding function can be used in its place.
Remark 4.30.
When choosing the initial virtual reference $\boldsymbol{\mathbf{v}}$ in the augmented system (24), it is worth noting that any $\boldsymbol{\mathbf{v}}_{0}\in\mathcal{V}$ such that $(\boldsymbol{\mathbf{x}}_{0},\boldsymbol{\mathbf{v}}_{0})\in\tilde{\mathcal{C}}$ is valid. A desirable choice is
$\boldsymbol{\mathbf{v}}_{0}=\underset{\boldsymbol{\mathbf{v}}\in\mathcal{V}}{\arg\min}~{}V(\boldsymbol{\mathbf{x}}_{0},\boldsymbol{\mathbf{v}}),$ (39)
because it tends to maximize the value of $\Delta$. If the above problem is infeasible, then $\boldsymbol{\mathbf{x}}_{0}$ is not a safe starting condition for the designed DSM.
Remark 4.31.
The proposed method effectively transforms the problem of finding a CBF into that of finding a Lyapunov function. The advantage in doing so is that the latter problem has a much richer history, which can be leveraged to synthesize Lyapunov-based DSMs using both analytical [30, 31, 32] and numerical [34] tools.
5 DSM-CBF Synthesis Approach Summary
In this manuscript, we have shown that dynamic safety margins (DSMs), key components in the explicit reference governor ERG, are (vector-valued) CBFs for an augmented system obtained by concatenating the state and reference vectors. This insight allows the direct use of approaches that compute DSMs to synthesize CBFs. Existing ERG works like [23] outline different approaches to construct DSMs. This manuscript studied the Lyapunov-based approach, which requires finding a reference-dependent Lyapunov function (33) for the prestabilized system and evaluating the safety (34) and stability (36) threshold values. These ingredients can then be combined to form a DSM. Once a DSM is obtained, a safe constrained control policy that minimally adjusts the nominal controller can be readily implemented by solving (31). The following examples illustrate using the Lyapunov-based approach to construct DSM-CBFs and implementing (31). Additional examples can be found in [39], where the approach presented in this paper is specialized to Euler–Lagrange systems.
6 Examples
In this section, we present two nonlinear examples to showcase the presented approach of designing CBFs using dynamic safety margins. For each example, we provide a comparison with candidate CBFs, backup CBFs, and the ERG. Simulations are computed in MATLAB and optimization problems are solved with MOSEK [40].
6.1 Anthill system
Consider the scalar system
$\dot{x}=(x^{2}-1)x+u.$ (40)
In the absence of inputs $u\!=\!0$, the system asymptotically tends to 00 if $|x_{0}|\!<\!1$ and diverges to infinity if $|x_{0}|>1$. Let $\mathcal{U}=[-u_{\max},u_{\max}]$ be the input constraint set, with $u_{\max}\!=\!\frac{4}{3\sqrt{3}}$.
Given the prestabilizing controller
$\pi(x,v)=-(v^{2}-1)v-3xv(x-v),$ (41)
with $v\in\mathbb{R}$, the prestabilized system is
$\dot{x}=f_{\pi}(x,v)=(x-v)^{3}-(x-v),$ (42)
which satisfies the equilibrium conditions $\bar{x}(v)=v$ and $\bar{u}(v)=\pi\big{(}\bar{x}(v),v\big{)}=v-v^{3}$ for all $v\in\mathbb{R}$. The equilibrium point $\bar{x}(v)$ is attractive if $|x_{0}-v|<1$. This can be proven using the Lyapunov function
$V(x,v)=\int_{v}^{x}\!\!\!\!-f_{\pi}(x,v)dx=\frac{1}{2}(x-v)^{2}-\frac{1}{4}(x-v)^{4}.$ (43)
This anthill-shaped function (see Fig. 2) is such that, given the $\mathscr{K}_{\infty}$ functions $\gamma_{1}(c)=\frac{1}{4}c^{2}$ and $\gamma_{2}(c)=\frac{1}{2}c^{2}$, $\forall(x,v)\in\tilde{\mathcal{D}}=\{(x,v)\in\mathbb{R}\times\mathbb{R}:x\in(v-1,v+1)\}$,
$\gamma_{1}(|x-v|)$ $\leq V(x,v)\leq\gamma_{2}(|x-v|),$
$\frac{\partial V}{\partial x}f_{\pi}(x,v)$ $=-f_{\pi}^{2}(x,v)\leq 0,$
The function $V(x,v)$ also satisfies Assumption 4. The stability threshold value (36) is
$\overline{\Gamma}(v)=V(v\pm 1,v)=\frac{1}{4}.$ (44)
To obtain the steady-state admissible set $\mathcal{V}\subset\mathbb{R}$, consider the input constraints at steady-state $|\bar{u}(v)|=|v-v^{3}|\leq u_{\max}$. Using Cardano’s formula to compute the roots of a 3-rd degree polynomial, we obtain $\mathcal{V}=\left[-v_{\max},v_{\max}\right]$, where
$v_{\max}=\frac{(2+\sqrt{3})^{\frac{1}{3}}+(2-\sqrt{3})^{\frac{1}{3}}}{\sqrt{3}}.$ (45)
The input constraints also induce state constraints for the prestabilized system. By setting $\pi(x,v)=\pm u_{\max}$ and solving for $x$, we obtain the reference-dependent state constraint set.
$\mathcal{X}_{v}=\begin{cases}\emptyset,\quad&v<-4/\sqrt{3},\\ \frac{1}{2}v+\left[\delta^{-}(v),-\delta^{-}(v)\right],\quad&v\in\left[-4/\sqrt{3},0\right),\\ \mathbb{R},\quad&v=0,\\ \frac{1}{2}v+\left[-\delta^{+}(v),\delta^{+}(v)\right],\quad&v\in\left(0,4/\sqrt{3}\right],\\ \emptyset,\quad&v>4/\sqrt{3},\end{cases}$
where
$\delta^{\pm}(v)=\frac{\sqrt{-3v^{4}+12v(v\pm u_{\max})}}{6v}.$ (46)
The resulting safety threshold value is
$\Gamma^{*}(v)\!=\!\begin{cases}-\infty,\;&v\in(-\infty,-1-v_{\max}],\\ -V\left(\frac{1}{2}v+\delta^{-}(v),v\right),\;&v\in(-1-v_{\max},-v_{\max}],\\ V\left(\frac{1}{2}v+\delta^{-}(v),v\right),\;&v\in(-v_{\max},1-v_{\max}),\\ \infty,\;&v\in[1-v_{\max},v_{\max}-1],\\ V\left(\frac{1}{2}v+\delta^{+}(v),v\right),\;&v\in(v_{\max}-1,v_{\max}),\\ -V\left(\frac{1}{2}v+\delta^{+}(v),v\right),\;&v\in[v_{\max},v_{\max}+1),\\ -\infty,\;&v\in[v_{\max}+1,\infty).\end{cases}$
While this function is discontinuous, it is bounded by the smooth function $\Gamma_{s}^{*}(v)=\max\left(\min\big{(}\Gamma^{*}(v),\frac{1}{4}\big{)},-\frac{1}{4}\right)$. Thus, by Theorem 4.25, $\Delta:\tilde{\mathcal{D}}\to\mathbb{R}^{2}$ defined as
$\Delta(x,v)=\begin{bmatrix}\Gamma_{s}^{*}(v)-V(x,v)\\ 0.99\overline{\Gamma}(v)-V(x,v)\end{bmatrix},$ (47)
is a DSM for $\pi$ and, by Theorem 3.6, a (vector-valued) CBF with respect to the augmented system.
Figure 2: Anthill-shaped Lyapunov function $V(x,v)$ at $v=0.5$, with the safety threshold value $\Gamma^{*}(v)$ and stability threshold value $\overline{\Gamma}(v)$.
We now consider the state constraints $\mathcal{X}=[-x_{\max},x_{\max}]$, with $x_{\max}=2$. Since $x_{\max}>v_{\max}$ and (40) is a first order system (no overshoot), these constraints have already been enforced by the restriction $v\in\mathcal{V}$ induced by $\tilde{\mathcal{C}}$. Thus, no further action is needed.
Remark 6.32.
It is interesting to note that, given $x_{\max}>v_{\max}$, the value of $x_{\max}$ plays no role in the DSM-based CBF. Indeed, the primary concern is to prevent instability by maintaining control authority over the system: any loss in control authority makes constraint violation inevitable, regardless of the specific value of $x_{\max}$.
Given the initial conditions $x_{0}=0.51$, $v_{0}=0.51$, the desired reference $r=1.5$, and the nominal control law $\kappa(x,r)=\pi(x,r)$, we consider the following constrained control strategies:
  1. (a)
    DSM-CBF: We use $\eta=0.01$, the class $\mathscr{K}_{\infty}$ function $\alpha:c\mapsto 1.8c$, and $\rho:(v,r)\mapsto r-v$.
  2. (b)
    ERG: We use $\rho_{\text{ERG}}:(v,r)\mapsto 100(r-v)$.
  3. (c)
    Candidate CBF: We use the functions $h_{1}(x)=x_{\max}-x$ and $h_{2}(x)=x_{\max}+x$ and tune their class $\mathscr{K}_{\infty}$ functions to illustrate the tradeoff between performance and feasibility that candidate CBFs suffer from.
  4. (d)
    Backup CBF: We consider the backup policy
    $\psi(x)=u_{\max}\left(\frac{1-e^{10x}}{1+e^{10x}}\right),$ (48)
    and formulate the Backup CBF
    $h(x)=\min_{\tau\in[0,10]}x_{\max}-\Phi_{f_{\psi}}(x,\tau),$ (49)
    where $\Phi_{f_{\psi}}(x_{0},\tau)$ is the solution of the backup dynamics $\dot{x}=f_{\psi}(x)\triangleq(x^{2}-1)x+\psi(x)$ at time $t=\tau$ with initial condition $x(0)=x_{0}$. The implementation is done following [21] with class $\mathscr{K}_{\infty}$ function $\alpha:c\mapsto 10c$.
The simulation results for each approach are shown in Fig. 3. Both the ERG and the proposed DSM-CBF successfully enforce safety and reach the closest safe point to the desired reference but, the DSM-CBF exhibits performance that more closely resembles the nominal behavior. With enough tuning, the candidate CBFs were able to also maintain safety for the fist 10 seconds, albeit with poor performance and no guarantee they will remain feasible afterwards. This is unsurprising since the candidate CBFs were constructed without any understanding of the inherent control authority constraint. The backup CBF also fails in this example. Although the backup policy $\psi$ is safe for the starting condition considered, it loses control authority as soon as $x>v_{\max}$, which causes it to become infeasible. Conversely, the DSM-based CBF provided a systematic approach that successfully identified the constraints induced by the input saturation.
Figure 3: Closed-loop behavior for the anthill system using each of the considered constrained control approaches. The dashed green and blue lines represent the evolution of the virtual reference $v(t)$ for the ERG and the DSM-CBF approaches, respectively. In addition, we show the performance of the candidate CBFs for different class $\mathscr{K}_{\infty}$ functions $\alpha_{1}=\alpha_{2}=\alpha:c\mapsto a_{i}c$, where $a_{i}\in\{7,0.25,0.15,0.1,0.07\}$. Larger $a_{i}$ corresponds to better input tracking, but also an earlier failure due to infeasibility, marked by a magenta ‘$\times$’ marker. The backup CBF is shown in orange and the ‘$\times$’ marker signals the point when the backup policy becomes unsafe, invalidating the approach.
6.2 Overhead Crane
Consider the dynamics of an overhead crane [41]
$M(\boldsymbol{\mathbf{q}})\ddot{\boldsymbol{\mathbf{q}}}+V_{m}(\boldsymbol{\mathbf{q}},\dot{\boldsymbol{\mathbf{q}}})\dot{\boldsymbol{\mathbf{q}}}+G(\boldsymbol{\mathbf{q}})=Bu,$ (50)
where the generalized coordinates $\boldsymbol{\mathbf{q}}=[x;~{}\theta]$ are the gantry position $x$ and the payload angle $\theta$, and
$M(\boldsymbol{\mathbf{q}})$ $=\begin{bmatrix}m_{c}+m_{p}&-m_{p}L\cos\theta\\ -m_{p}L\cos\theta&m_{p}L^{2}\end{bmatrix},\quad B=\begin{bmatrix}1\\ 0\end{bmatrix},$
$V_{m}(\boldsymbol{\mathbf{q}},\dot{\boldsymbol{\mathbf{q}}})$ $=\begin{bmatrix}0&m_{p}L\dot{\theta}\sin\theta\\ 0&0\end{bmatrix},\quad G(\boldsymbol{\mathbf{q}})=\begin{bmatrix}0\\ m_{p}gL\sin\theta\end{bmatrix},$
where $m_{c},m_{p}>0$ represent the gantry and payload masses, respectively, $L>0$ is the length of the rod connecting the crane and payload, and $g=9.81$ is the acceleration of gravity (see Fig. 4).
Figure 4: Overhead crane system from [41].
It can be shown that, letting $\boldsymbol{\mathbf{x}}=[\boldsymbol{\mathbf{q}};~{}\dot{\boldsymbol{\mathbf{q}}}]$, the system is control-affine and can be written as in (1). We consider the prestabilizing PD control law given in [41]
$\pi(\boldsymbol{\mathbf{x}},v)=-k_{p}(x-v)-k_{d}\dot{x},$ (51)
where $k_{p}>0$, $k_{d}>0$ are the proportional and derivative gains. The equilibrium mapping is $\bar{x}:v\mapsto[v~{};0~{};0~{};0]$ and the reference-dependent Lyapunov function is
$V(\boldsymbol{\mathbf{x}},v)=\frac{1}{2}\dot{\boldsymbol{\mathbf{q}}}^{\top}M(\boldsymbol{\mathbf{q}})\dot{\boldsymbol{\mathbf{q}}}+m_{p}gL(1-\cos\theta)+\frac{1}{2}k_{p}(x-v)^{2},$
which satisfies Definition 4.21 with $\mathcal{D}_{v}=\mathbb{R}\times(-\pi,\pi)\times\mathbb{R}^{2}$ and also satisfies Assumption 4. We now define four types of constraints and, for each one, we find the associated safety threshold value or a continuously differentiable bounding function. For details on the construction of each $\Gamma^{*}_{i}(v)$, the reader is referred to [30, 39].
  1. 1.
    Position constraints: $x_{\text{min}}\leq x\leq x_{\text{max}}$. For each of the two bounds, we take
    $\Gamma_{1}^{*}(v)$ $=\frac{1}{2}k_{p}(v-x_{\text{min}})|v-x_{\text{min}}|,$
    $\Gamma_{2}^{*}(v)$ $=\frac{1}{2}k_{p}(x_{\text{max}}-v)|x_{\text{max}}-v|.$
  2. 2.
    Input constraints: $|u|\leq u_{\text{max}}$, with $u_{\text{max}}>0$. Solving (34) yields the constant function
    $\Gamma^{*}_{3}(v)=\frac{m_{c}u_{\text{max}}^{2}}{2(m_{c}k_{p}+k_{d}^{2})}.$
  3. 3.
    Angle constraints: $|\theta|\leq\theta_{\text{max}}$, with $\theta_{\text{max}}\in(0,\pi/2)$. Solving (34) yields the constant function
    $\Gamma_{4}^{*}(v)=m_{p}gL(1-\cos\theta_{\text{max}}).$
  4. 4.
    Payload constraints: $x+L\sin\theta\leq p_{\text{max}}$, with $p_{\text{max}}\geq x_{\text{max}}$. For this constraint, we were unable to find an analytical expression for $\Gamma_{5}^{*}(v)$ associated to $V$. However, as detailed in [30, Proposition 4], we can define
    $\underline{V}(\boldsymbol{\mathbf{x}},v)=\frac{1}{2}\dot{\boldsymbol{\mathbf{q}}}^{\top}M(\boldsymbol{\mathbf{x}})\dot{\boldsymbol{\mathbf{q}}}+\frac{4}{\pi^{2}}m_{p}gL\theta^{2}+\frac{1}{2}k_{p}(x-v)^{2},$
    which satisfies $\underline{V}(\boldsymbol{\mathbf{x}},v)\leq V(\boldsymbol{\mathbf{x}},v)$, $\forall\theta\in(-\pi/2,\pi/2)$. Moreover, since $x_{\text{max}}\leq p_{\text{max}}$, we can perform a planar embedding of the payload constraint
    $x+L\theta\leq p_{\text{max}}\quad\implies\quad x+L\sin\theta\leq p_{\text{max}},$ (52)
    which holds $\forall x\leq x_{\text{max}}$ and $\forall\theta\in[-\pi/2,\pi/2]$. With these approximations, the following continuously differentiable function is a bound for $\Gamma^{*}_{5}(v)$ and can be used in its stead:
    $\Gamma_{5}(v)=\frac{4k_{p}m_{p}g}{8m_{p}g+Lk_{p}\pi^{2}}(p_{\text{max}}-v)|p_{\text{max}}-v|.$
The stability threshold value is
$\overline{\Gamma}(v)=2m_{p}gL.$
Using Theorem 4.25, we conclude that
$\Delta(\boldsymbol{\mathbf{x}},v)=\begin{bmatrix}\Gamma^{*}_{1}(v)-V(\boldsymbol{\mathbf{x}},v)\\ \Gamma^{*}_{2}(v)-V(\boldsymbol{\mathbf{x}},v)\\ \Gamma^{*}_{3}(v)-V(\boldsymbol{\mathbf{x}},v)\\ \Gamma^{*}_{4}(v)-V(\boldsymbol{\mathbf{x}},v)\\ \Gamma_{5}(v)-V(\boldsymbol{\mathbf{x}},v)\\ 0.99\overline{\Gamma}(v)-V(\boldsymbol{\mathbf{x}},v)\end{bmatrix},$
is a DSM for $\pi$ and, by Theorem 3.6, a (vector-valued) CBF with respect to the augmented system. Note that the stability threshold value satisfies $\Gamma^{*}_{4}(v)<\overline{\Gamma}(v)$. As such, when the angle constraints are enforced, there is no need to include $\overline{\Gamma}(v)$ in the definition of $\Delta$ above.
For all simulations, we use linear class $\mathscr{K}_{\infty}$ functions given by $\alpha_{i}:c\mapsto\alpha_{i}c$ where $\alpha_{i}>0$ represents both the function and the scalar gain despite the slight abuse of notation. We choose $\alpha_{1}=\alpha_{2}=10$, $\alpha_{3}=20$, $\alpha_{4}=80$, $\alpha_{5}=2.2$ and $\alpha_{6}=80$. The navigation field we use is $\rho(v,r)=r-v$.
We compare with the following approaches.
  1. (a)
    Nominal: The nominal controller is a PD control law $\kappa(\boldsymbol{\mathbf{x}},r)=-k_{p,\kappa}(x-r)-k_{d,\kappa}\dot{x}$, where the gains $k_{p,\kappa}>0$ and $k_{d,\kappa}>0$ can differ from the ones in (51) and $r\in\mathbb{R}$ is the target reference.
  2. (b)
    ERG: The explicit reference governor uses the same dynamic safety margins as our approach, but changing $k_{p}\to k_{p,\kappa}$ and $k_{d}\to k_{d,\kappa}$ in all the expressions. The navigation field is $\rho_{\text{ERG}}(v,r)=1000(r-v)$. The ERG-based, closed-loop system is
    $\begin{bmatrix}\dot{\boldsymbol{\mathbf{x}}}\\ \dot{v}\end{bmatrix}=\begin{bmatrix}f(\boldsymbol{\mathbf{x}})+g(\boldsymbol{\mathbf{x}})\kappa(\boldsymbol{\mathbf{x}},v)\\ \min_{i}\Delta_{i}(\boldsymbol{\mathbf{x}},v)\rho_{\text{ERG}}(v,r)\end{bmatrix}.$ (53)
  3. (c)
    Candidate CBF: For each constraint, we design a candidate CBF.
    1. 1)
      Position constraints: $h_{1}(\boldsymbol{\mathbf{x}})=\gamma_{1}(x-x_{\text{min}})+\dot{x}$ for $x_{\text{min}}\leq x$ and $h_{2}(\boldsymbol{\mathbf{x}})=\gamma_{2}(x_{\text{max}}-x)-\dot{x}$ for $x\leq x_{\text{max}}$.
    2. 2)
      Input constraints are added as upper and lower bounds for the optimization variable $-u_{\text{max}}\leq u\leq u_{\text{max}}$.
    3. 3)
      Angle constraints: $h_{3}(\boldsymbol{\mathbf{x}})=\gamma_{3}(\theta_{\text{max}}+\theta)+\dot{\theta}$ for $-\theta_{\text{max}}\leq\theta$ and $h_{4}(\boldsymbol{\mathbf{x}})=\gamma_{4}(\theta_{\text{max}}-\theta)-\dot{\theta}$ for $\theta\leq\theta_{\text{max}}$.
    4. 4)
      Payload constraints: $h_{5}(\boldsymbol{\mathbf{x}})=\gamma_{5}(p_{\text{max}}-x-L\sin\theta)-\dot{x}-L\dot{\theta}\cos\theta$ for $x+L\sin\theta\leq p_{\text{max}}$.
    Where $\gamma_{i}$ are positive constants. For all simulations, we use $\gamma_{i}=6$ and $\tilde{\alpha}_{i}=8,\forall i\in\{1,\ldots,4\}$, and $\gamma_{5}=4$, $\tilde{\alpha}_{5}=3$, where $\tilde{\alpha}_{i}:c\mapsto\tilde{\alpha}_{i}c$ is the class $\mathcal{K}_{\infty}$ function associated with $h_{i}$. It must be noted that all $h_{i}$ described above are candidate CBFs, meaning that they may not satisfy Definition 2 or have the control-sharing property given in Definition 2.4.
  4. (d)
    Backup CBF: We consider the backup policy
    $\psi(\boldsymbol{\mathbf{x}})=-u_{\max}.$ (54)
    For each constraint, we design a backup CBF and implement it according to [21]. For instance, the backup CBF associated to the payload constraint is
    $h(\boldsymbol{\mathbf{x}})\!=\!\min_{\tau\in[0,10]}p_{\max}\!-\!\boldsymbol{\mathbf{e}}_{1}^{\top}\Phi_{f_{\psi}}(\boldsymbol{\mathbf{x}},\tau)\!-\!L\sin\left(\boldsymbol{\mathbf{e}}_{2}^{\top}\Phi_{f_{\psi}}(\boldsymbol{\mathbf{x}},\tau)\right),$
    where $\boldsymbol{\mathbf{e}}_{i}$ is the $i$-th column of $I_{4}$ and $\Phi_{f_{\psi}}(\boldsymbol{\mathbf{x}}_{0},\tau)$ is the solution of the backup dynamics $f_{\psi}$ at time $t=\tau$ with initial condition $\boldsymbol{\mathbf{x}}(0)=\boldsymbol{\mathbf{x}}_{0}$. All class $\mathscr{K}_{\infty}$ functions are the same $\alpha(c)=300c$.
Table 1: Overhead Crane Simulation Parameters
$m_{c}$ (kg) $1$ $k_{p}$ $1$ $x_{\text{min}}$ (m) $-1.1$ $p_{\text{max}}$ (m) $1.1$
$m_{p}$ (kg) $0.5$ $k_{d}$ 0.1 $x_{\text{max}}$ (m) $1.1$ $\boldsymbol{\mathbf{x}}(0)$ (-) 00
$L$ (m) $0.7$ $k_{p,\kappa}$ $10$ $u_{\text{max}}$ (N) $4$ $v(0)$ (m) $0.1$
$\eta$ (-) 0.02 $k_{d,\kappa}$ $4$ $\theta_{\text{max}}$ (deg) $10$ $r$ (m) $1$
General parameters for the example are provided in Table 1. For the first simulation, shown in Fig. 5, we remove the angle constraint. In this case, all constrained control approaches achieve the safety objective. In terms of performance, the backup and candidate CBF approaches track the nominal behavior better than the proposed DSM-CBF approach, whereas the ERG exhibits the slowest response. The second simulation, shown in Fig. 6, includes the angle constraint $|\theta|\leq\theta_{\text{max}}$. The ERG and our proposed DSM-CBF approaches successfully enforce the constraints, with the proposed method achieving better performance. However, the candidate CBF-based controller becomes infeasible despite our best efforts to tune $\gamma_{i}$ and $\tilde{\alpha}_{i}$. This is unsurprising since candidate CBFs do not guarantee persistent feasibility. The backup policy $\psi$ is unsafe under the introduced angle constraints. This highlights that finding a suitable backup policy becomes more challenging as the number of constraints increases.
Figure 5: Closed-loop behavior in the absence of angle constraints. Even though the nominal controller violates the constraints, the ERG, candidate CBF, backup CBF and proposed DSM-CBF approaches all successfully enforce the safety requirements. The dashed blue and green lines are the virtual reference $v$ of the DSM-CBF and ERG approaches, respectively. Note that $v(t)$ may decrease in our approach, which is not possible in the ERG. This feature enlarges the feasible region of the CBF-based QP and allows the safe input $\beta(x,v)$ to be closer to the nominal input $\pi(x,r)$.
Figure 6: Closed-loop behavior in the presence of angle constraints. The ‘$\times$’ marks the instant in which the candidate and backup CBF policies become infeasible. The failure in the candidate CBF is likely due to the unavoidable backswing experienced by the payload angle $\theta(t)$. The failure in the backup CBF is due to the backup policy being unsafe with respect to the angle constraints. Conversely, the ERG and proposed DSM-CBF approaches successfully enforce the safety requirements.
7 Conclusion
This paper showed that dynamic safety margins are control barrier functions for an augmented system obtained by concatenating the state and reference vectors. The work presents an optimization-based safe policy and it is shown that it is persistently feasible. Then, a Lyapunov-based approach to construct DSMs was presented in detail and its performance against other common approaches to design CBFs, as well as against the ERG was compared in two nonlinear systems.
Future work includes continuing to explore connections between the CBF and reference governor literatures, adapt other types of DSMs to construct CBFs, and robustifying the presented results to account for bounded disturbances.